### v3

# Longhorn: State Space Models are Amortized Online Learners

Abstract
The most fundamental capability of modern AI methods such as Large Language Models (LLMs) is the ability to predict the next token in a long sequence of tokens, known as “sequence modeling.” Although the Transformers model is the current dominant approach to sequence modeling, its quadratic computational cost with respect to sequence length is a significant drawback. State-space models (SSMs) offer a promising alternative due to their linear decoding efficiency and high parallelizability during training. However, existing SSMs often rely on seemingly ad hoc linear recurrence designs. In this work, we explore SSM design through the lens of online learning, conceptualizing SSMs as meta-modules for specific online learning problems. This approach links SSM design to formulating precise online learning objectives, with state transition rules derived from optimizing these objectives. Based on this insight, we introduce a novel deep SSM architecture based on the implicit update for optimizing an online regression objective. Our experimental results show that our models outperform state-of-the-art SSMs, including the Mamba model, on standard sequence modeling benchmarks and language modeling tasks. [^1] Code is available at https://github.com/Cranial-XIX/longhorn.git.
Figure 1: (left) The average perplexity across eight downstream datasets for GLA, Mamba and Longhorn (1.3B model) over seen tokens on SlimPajama. Longhorn leads to a 1.8x speed up in sampling efficiency. (right) Longhorn, pretrained with 2048 context length, successfully extrapolates to longer context (up to 16x of the trained context).
1 Introduction
The Transformer model has become the go-to architecture for sequence modeling in deep learning (Vaswani et al., 2017). However, its utility is constrained by the quadratic growth in training and decoding costs with increasing sequence length. Despite various optimizations such as efficient decoding (Chen et al., 2023; Kuperman & Dyke, 2011), KV-cache compression (DeepSeek-AI & Dai, 2024), and memory efficient implementation (Dao et al., 2022), scaling Transformers for autonomous and continual use with an infinite (or very long) context window remains a challenge.
Recent advancements in linear attention models (Katharopoulos et al., 2020) and state-space models (SSMs)(Gu et al., 2021) have demonstrated their potential. These models are specialized recurrent neural networks capable of efficiently computing outputs in parallel when input tokens are provided simultaneously during training, thus avoiding the inefficiencies of traditional backpropagation through time. During inference, the recurrent form is employed, resulting in linear decoding efficiency. Initially, these models underperformed compared to Transformers. However, recent SSMs (Gu & Dao, 2023; Yang et al., 2023; Peng et al., 2024; De et al., 2024; Beck et al., 2024) have achieved performance parity with Transformers in language modeling tasks. Despite extensive research into various design aspects of SSMs, a guiding principle for designing SSMs remains elusive.
In this work we propose one potential principle. We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history into a memory state which is then used by later layers in the network for sequence modeling. From this perspective:
The recurrent form of SSMs can be viewed as solving an online learning problem.
Figure 2: (left) Most existing sequence models consist of channel and sequence mixing layers. The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $s_{t}$, which is then passed to later layers for sequence modeling. (middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows a certain dynamics that optimizes an online learning objective. The recurrent update of $s$ can then be derived by solving the online objective in closed-form or by conducting a gradient step. (right) Longhorn’s dynamics is derived from solving an online regression problem, where the goal is to recover $x\in\mathbb{R}$ (if $x\in\mathbb{R}^{d}$ we can run $d$ SSMs in parallel) based on a hint $k\in\mathbb{R}^{m}$ from a state vector $s\in\mathbb{R}^{d}$. The closed-form solution to the online regression problem corresponds to the implicit online learning update, where $\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$. Please see the details in Section 3.
As a result, we can draw inspiration from online learning and confine the design choices of SSMs to reflect those learning dynamics that solve specific online prediction problems. The hope is that by selecting the right objective, one can save parameters / computation without sacrificing performance. In addition, the proposed online learning perspective might lead to better understanding of what SSM layers are doing in a large model. Specifically, we view the recurrent update (a.k.a., the state-transition dynamics) of an SSM model as either a gradient descent step or a closed-form solution to an online learning objective. We provide the objectives for some existing SSMs in Table 1.
Based on this insight, we propose a simple yet effective architecture (Longhorn), derived from the implicit closed-form update of an online associative recall problem. The closed-form update naturally leads to a stable recurrent form without a manually designed gating mechanism. So Longhorn does not need a separately parameterized forget gate, which saves parameters when the state size is large. We demonstrate that Longhorn performs comparably to or better than state-of-the-art SSMs like Mamba (Gu & Dao, 2023) on both synthetic and large-scale sequence modeling tasks. In particular, Longhorn outperforms Mamba at the size of 1.3B-parameter when trained on 100B tokens from the SlimPajama dataset (Soboleva et al., 2023). To summarize, our contributions are:
1) Theoretical Framework: We propose a novel framework that views SSM’s recurrent update as solving online learning objectives. As a result, the design of SSMs reduces to the design of the online learning objectives. In particular, we introduce a novel, simple, and effective SSM, named Longhorn, that explicitly solves an online associative recall problem. Longhorn’s recurrent update is obtained by the closed-form solution to the online learning objective. Consequently, Longhorn does not require a separately parameterized forget gate that appears in most existing SSMs.
2) Empirical Results: Longhorn demonstrates better performance than existing SSMs including Mamba, across both synthetic associative recall tasks and the large-scale language modeling task. Moreover, it achieves 1.8x enhancement in sampling efficiency compared to Mamba (See Figure 1 (left)). Longhorn’s training speed is as fast as Mamba, as we only replace the SSM module in the Mamba architecture with Longhorn’s recurrence. So it serves as a drop-in replacement for Mamba. Lastly, Longhorn, trained with 2048 context length can extrapolate to 32K context length at inference time without much perplexity drop (See Figure 1 (right)).
Notation
[…]
Algorithm 1 Longhorn’s Single-layer SSM Recurrence (Inference Time)
1:  Parameters: $W_{q}\in\mathbb{R}^{m\times d},W_{k}\in\mathbb{R}^{m\times d},W_{\beta}\in\mathbb{R}^{d\times d}$, where $W_{\beta}$ can be low-rank, horizon $T$.
2:  Initialize the memory state $S_{0}\leftarrow 0^{d\times m}$.
3:  for $t\in\{1,\dots,T\}$ do
4:     1) Receive input $x_{t}\in\mathbb{R}^{d}$.
5:     2) Compute the query $q_{t}$, key $k_{t}$ and $\beta_{t}$ (as in objective $||s-s_{t-1}||^{2}+\beta_{t}||s^{\top}k_{t}-x_{t}||^{2}$):
$q_{t}=W_{q}x_{t}\in\mathbb{R}^{m},$ $k_{t}=W_{k}x_{t}\in\mathbb{R}^{m},$ $\beta_{t}=\text{Sigmoid}(W_{\beta}x_{t})\in(0,1)^{d}.$
6:     3) Update the memory state $S_{t}\in\mathbb{R}^{d\times m}$ via
$S_{t}=\big{(}1-\Delta_{t}\otimes k_{t}^{\odot 2}\big{)}\odot S_{t-1}+\big{(}\Delta_{t}\odot x_{t}\big{)}\otimes k_{t},$
where $\Delta_{t}$ is the step size:
$\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})\in(0,1)^{d}.$
7:     4) Compute the output $o_{t}=S_{t}q_{t}\in\mathbb{R}^{d}$.
8:  end for
9:  Note: $\odot$ elementwise product and $\otimes$ is outer product. $x_{t}$ in practice is preprocessed through a linear projection followed by a Conv1d operation as in Mamba (Gu & Dao, 2023).
2 Background
In this section, we provide a brief introduction to contemporary deep state space models (deep SSMs).
Modern large language models are sequence-to-sequence models consisting of a stack of layers $\bm{y}=\Phi_{L}\circ\cdots\circ\Phi_{1}(\bm{x})$ that sequentially processes an input sequence $\bm{x}=\{x_{t}\}_{t=1}^{T}$, where $T$ is the context length. Specifically, transformers consist of alternative stacks of self-attention (SA) and multi-layer perceptron (MLP) layers that conduct mixing (i.e., information aggregation) on the sequence and channel dimensions, respectively.
Deep SSMs replace the SA layers with SSM layers. Some variants of SSM models leave the MLP layers unchanged (Sun et al., 2023; Yang et al., 2023; De et al., 2024), while others fuse the SSM layer and the MLP layer into a single unified module (Gu & Dao, 2023). But in both cases, the sequence mixing is done by the SSM module and the channel mixing is done by the channel-wise MLP. Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of a SSM block for sequence mixing (red), and a MLP block for channel mixing (blue).
SSM: General Form
The SSM block (in red) plays the crucial role of sequence mixing. It works by iteratively updating a memory state matrix $S_{t}\in\mathbb{R}^{d\times m}$ with a linear recurrence:
$S_{t}=A(x_{t})*S_{t-1}+B(x_{t}),$ $\forall t\in\{1,\ldots,T\},$ $S_{0}=0,$ (1)
where $x_{t}$ is the input at time $t$, $S_{t}$ is the model’s state, $A_{t},B_{t}\colon\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are some functions and * is a multiplication operation of choice, such as Hadamard product or matrix product.
Given the state $S_{t}$, SSMs often give the output token at the next layer via a gated linear unit (GLU) (Dauphin et al., 2017):
$y_{t}=\mathtt{Readout}(S_{t},x_{t})=W_{1}\big{(}o_{t}\odot\sigma(W_{2}x_{t})\big{)},$ $o_{t}=C(x_{t})S_{t},$
where we first get $o_{t}$ via a state-dependent linear projection on $S_{t}$, which is then fed into a subsequent channel mixing gated linear unit (blue in Figure 3), where $\sigma(\cdot)$ is a non-linear activation function.
Figure 3: Mamba Block
A key feature of this design in Equation 1 is that the update of $S_{t}$ is a linear recurrence relation. That is, $S_{t}$ is a linear function of $S_{t-1}$. Crucially, this allows us to express all $S_{t}$ in an explicit form that can be calculated in parallel: when all $\bm{x}=\{x_{t}\}_{t}$ are available as in the training phase, $\{S_{t}\}_{t}$ can be written into
$S_{t}=\sum_{t^{\prime}\leq t}(\overline{A}_{t^{\prime}\to t})B(x_{t^{\prime}}),~{}~{}~{}~{}~{}~{}\text{where}~{}~{}~{}~{}~{}~{}~{}\overline{A}_{t^{\prime}\to t}=\prod_{t^{\prime}<\tau\leq t}A(x_{\tau}).$ (2)
Here $\prod$ denotes the product induced by multiplication operator *. The resulting cumulative product $\overline{A}_{t^{\prime}\to t}$ can be implemented efficiently in parallel with the prefix scan algorithm (Harris et al., 2007), which only requires a sublinear complexity in terms of sequence length (e.g., $\mathcal{O}(\log t)$). From now on, we will abbreviate $A(x_{t})$ and $B(x_{t})$ as $A_{t}$ and $B_{t}$, respectively.
Designs of ($A_{t}$, $B_{t}$, *)
Existing variants of SSMs mainly differs in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence. A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed be $m$ times of the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence. This makes the architecture design of $A_{t},B_{t}$, both mapping $\mathbb{R}^{d}$ to $\mathbb{R}^{d\times m}$ challenging. A naive linear lay would result in $d\times d\times m$ weights, which is prohibitively large. This makes it necessary to impose certain low dimensional structure in $A_{t},B_{t}$, which is what differentiate the existing designs of SSM. In the following, we provide some examples in the form of Equation 1 for some existing SSM models.
Example 2.1 (Linear Attention Variants).
Linear Attention (LA) (Katharopoulos et al., 2020), Retention Network (RetNet) (Sun et al., 2023), and Gated Linear Attention (GLA) (Yang et al., 2023) all assume $A_{t},B_{t}$ yield rank-1 (or even constant) outputs:
$S_{t}=A_{t}\odot S_{t-1}+v(x_{t})\otimes k(x_{t}),~{}~{}~{}~{}\text{with}~{}~{}\begin{cases}A_{t}=1&\quad\text{(LA)}\\ A_{t}=c\in[0,1]&\quad\text{(RetNet)}\\ A_{t}=1\otimes\alpha(x_{t})&\quad\text{(GLA)}\end{cases},$
where $S_{t}\in\mathbb{R}^{d\times m}$, $v(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ are linear mappings of $x_{t}$, and $\otimes$ denote the outer product.
In practice, one can choose to use $h$ heads as in the multi-head attention to save some computation, where both the $m$ and $d$ dimension are divided into $h$ groups and each group performs its own LA variant. The outer product becomes $\mathcal{O}(h*m/h*d/h=md/h)$. But then the effective size of $S_{t}$ also shrinks to $md/h$.
Example 2.2 (Mamba (Gu & Dao, 2023)).
The Mamba architecture is derived by discretizing a continuous linear dynamics. Its discretized update is:
$\begin{split}S_{t}&=A_{t}\odot S_{t-1}+B_{t},~{}~{}~{}~{}\text{where}\\ A_{t}&=\exp(A\odot(\Delta(x_{t})\otimes 1)),\qquad B_{t}=(\Delta(x_{t})\odot x_{t})\otimes k(x_{t}).\end{split}$ (3)
where $S_{t}\in\mathbb{R}^{d\times m}$ with $m=16$ by default, $\Delta(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ linear mappings of $x_{t}$, and $A\in\mathbb{R}^{d\times m}$ is a data independent (not depending on $x_{t}$ trainable weight matrix.
We can see that in Mambda both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
In practice, Mamba does not use multiple heads as in linear attention variants. Perhaps the main reason is that given a fixed $m$ and $d$, the largest memory state will be with $h=1$ (as the effective size of $S_{t}$ is $md/h$). In addition, Mamba’s output is $o_{t}=C(x_{t})S_{t}+D_{t}\odot x_{t}$, which has an additional residual part $D_{t}\odot x_{t}$.
Example 2.3 (Griffin (De et al., 2024)).
In Mamba and the linear attention variants, the outer product serves a critical rule of lifting vectors to matrices. The recent Griffin architecture abandons the outer product and performs pure elementwise product:
$s_{t}=a(x_{t})\odot s_{t-1}+\sqrt{1-a(x_{t})}\odot i(x_{t})\odot x_{t},$
where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$. This yields smaller memory states, but in practice Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
Example 2.4 (RWKV (Peng et al., 2023)).
The original RWKV also performs elementwise recurrence. It maintains a state of ratio form $s_{t}=u_{t}/z_{t}$, where $u_{t},z_{t}$ are updated separately by two SSMs:
$s_{t}$ $=u_{t}/z_{t}$
$u_{t}$ $=\exp(-w)\cdot u_{t-1}+\exp(k(x_{t}))\odot v(x_{t}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}z_{t}=\exp(-w)\cdot z_{t-1}+\exp(k(x_{t})),$
where all the vectors are of size $\mathbb{R}^{d}$, and $w>0$ is a trainable weight for controlling the forgetting. In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with outer product, which makes it more similar to a LA variant.
Example 2.5 (HGRN2 (Qin et al., 2024a)).
The Gated Linear RNNs with State Expansion (HGRN2) model is represented with the following recurrence:
$S_{t}=(1\otimes f(x_{t}))\odot S_{t-1}+i(x_{t})\otimes(1-f(x_{t})).$
Here, $f(x_{t})\in[0,1]$ is the forget gate, $(1-f(x_{t}))$ is the input gate, and $i(x_{t})$ is the input vector. HGRN2 thus resembles an RNN.
3 An Online Learning Perspective For Sequence Mixing
As demonstrated in the previous section, the design of a state-space model (SSM) depends on the specific selection of $(A_{t},B_{t},*)$, a process that is both intricate and somewhat artisanal. In this section, we propose to streamline SSM design through a online learning perspective. The main idea is to treat the SSM layers as learning modules that learn to compress information along the sequence dimension. From this perspective, the SSM layers are learning to learn, such that at the inference time, these layers are still learning (compressing) new information online.
We begin with an overview of online learning and subsequently demonstrate how SSM can be framed as an online learning problem. Finally, we present a straightforward architecture based on the closed-form solution of the implicit online learning algorithm.
3.1 SSM as Online Learning
We advocate to view the recurrence of SSM as solving an online learning problem. In online learning, the agent (the algorithm) picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$. The goal is to minimize
$\min_{\{s_{t}\}}\sum_{t}\ell_{t}(s_{t}).$ (4)
For instance, consider online linear prediction, where at each step the agent is given an input-label pair $(x_{t},y_{t})$ and $\ell_{t}(s_{t})=\frac{1}{2}||s_{t}^{\top}x_{t}-y_{t}||^{2}$ is the $\ell_{2}$ regression loss, then the problem becomes an online regression problem, and the goal is to successfully predict $y_{t}$ given $x_{t}$ at future time steps.
Online convex optimization (OCP) yields a principled approach to solving Equation 4 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasiticity” (Mermillod et al., 2013). Formally, an online convex programming algorithm update $s_{t}$ by solving a regularized cost function:
$s_{t}=\operatorname*{arg\,min}_{s}L_{t}(s),~{}~{}~{}~{}~{}~{}~{}L_{t}(s)=\underbrace{D_{\phi}(s,s_{t-1})}_{\text{stability}}+\underbrace{\beta_{t}\ell_{t}(s)}_{\text{plasticity}},$ (5)
where $\beta_{t}\in\mathbb{R}^{+}$ and $D_{\phi}$ is a discrepancy measure, often a Bregman divergence induced by the convex function $\phi$ (e.g., when $\phi(x)=\frac{1}{2}\left\lVert x\right\rVert^{2}$, $D_{\phi}(s,s_{t-1})=\frac{1}{2}||s-s_{t-1}||^{2}$). Here the first term ensures the updated $s$ will be close to the previous $s_{t-1}$, so the agent suffers less from catastrophic forgetting, while the second term ensures the agent is incorporating new knowledge from minimizing the new loss $\ell_{t}(s)$. Hence, $\beta_{t}$ controls the trade-off between stability and plasiticity.
3.2 SSM as Online Learning and The Longhorn Architecture
Under the online learning framework, the design of SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 5. This provides a unified framework for the existing SSM variants. We summarize in Table 1 the online learning interpolation of a number of existing SSM architectures.
In this work, we explore a highly simplified and natural design called Longhorn guided the online principle (see the last row of Table 1). In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is higher dimensional). Then, we propose to use the following OCP objective:
$s_{t}=\operatorname*{arg\,min}_{s}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
Here, $k_{t}(x_{t})\in\mathbb{R}^{d}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer. $\beta_{t}(x_{t})$ controls how much new information about $x_{t}$ we want the model to incorporate. For instance, $\beta_{t}(x_{t})=0$ implies $s_{t}=s_{t-1}$, while a large $\beta_{t}(x_{t})$ implies the model sacrifices stability for plasticity. For simplicity, we will refer to them as $k_{t},\beta_{t}$ from now on. From a high-level, Equation 6 is solving an online prediction problem, also known as the associative memory problem (a supervised learning problem), where the goal is to memorize all $(k_{t},x_{t})$ pairs (learn a mapping from $k_{t}$ to $x_{t}$) such that given a key (input) $k$ the model can retrieve (predict) its corresponding value (label) $x$.
Fortunately, this simple objective gives a closed form solution for $s_{t}$ (a.k.a., the implicit online learning method (Kulis & Bartlett, 2010)).
Theorem 3.1.
The closed form solution for $s_{t}$ for objective in Equation 6 is
$s_{t}=(I-\Delta_{t}k_{t}k_{t}^{\top})s_{t-1}+\Delta_{t}k_{t}x_{t},~{}~{}~{}\text{where}~{}\Delta_{t}=\frac{\beta_{t}}{1+\beta_{t}k_{t}^{\top}k_{t}}\in[0,\infty).$ (7)
Proof.
As the objective in Equation 6 is in a quadratic form with respect to $s$, there is a unique minimum:
$\nabla L_{t}(s)=0$ $\Longleftrightarrow(s-s_{t-1})+\beta_{t}(s^{\top}k_{t}-x_{t})k_{t}=0$
$\Longleftrightarrow(I+\beta_{t}k_{t}k_{t}^{\top})s=s_{t-1}+\beta_{t}k_{t}x_{t}$
$\underbrace{\Longleftrightarrow}_{(3)}s=\bigg{(}I-\frac{\beta_{t}}{1+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}\beta_{t}k_{t}x_{t}$
$\Longleftrightarrow\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\frac{(I+\beta_{t}k_{t}^{\top}k_{t}-\beta_{t}k_{t}k_{t}^{\top})\beta_{t}k_{t}x_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}$
$\underbrace{\Longleftrightarrow}_{(5)}\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\frac{\beta_{t}k_{t}x_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}$
(3) is derived from the fact that $(I+\beta_{t}k_{t}k_{t}^{\top})^{-1}=(I-\frac{\beta_{t}k_{t}k_{t}^{\top}}{I+\beta k_{t}^{\top}k_{t}})$ by the Sherman–Morrison formula. (5) is derived by noticing that $k_{t}^{\top}k_{t}k_{t}x_{t}-k_{t}k_{t}^{\top}k_{t}x_{t}=0$. ∎
As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting parallel scan. As a result, in practice, we use the diagonal approximation $1-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$. The final SSM recurrence therefore becomes:
$s_{t}=\underbrace{(1-\Delta_{t}k_{t}^{\odot 2})}_{A_{t}(x_{t})}\odot s_{t-1}+\underbrace{\Delta_{t}k_{t}x_{t}}_{B_{t}(x_{t})}.$ (8)
When $x_{t}\in\mathbb{R}^{d}$ is a higher dimensional vector, following the Mamba design, we use $d$ parallel SSMs. As a result, $x_{t},\beta_{t},\Delta_{t}\in\mathbb{R}^{d},k_{t}\in\mathbb{R}^{m},S_{t}\in\mathbb{R}^{d\times m}$, Longhorn’s update becomes:
$S_{t}=A_{t}\odot S_{t-1}+B_{t},~{}~{}\text{where}~{}~{}A_{t}=(1-\Delta_{t}\otimes k_{t}^{\odot 2}),~{}~{}B_{t}=(\Delta_{t}\odot x_{t})\otimes k_{t}.$ (9)
The final architecture of Longhorn follows Mamba strictly (Figure 3), and we only replace the SSM block with the Longhorn’s recurrence. We also provide an efficient CUDA kernel for it. The full inference time algorithm is provided in Algorithm 1. One can compare Equation 9 to Equation 3 and other SSM in Section 2. Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forget gate is naturally derived from the key vector, i.e., $(1-k^{\odot 2}\otimes\Delta_{t})$.
Advantages of Longhorn: 1. While we can derive the learning objective for some of the existing SSMs, Longhorn is the first SSM designed for explicitly solving an online regression problem. 2. Longhorn does not require a specific forget gate (e.g., $\alpha_{t}$ in GLA or $A$ matrix in Mamba). The forgetting is naturally linked to the key vector $k_{t}$ through the derivation. This saves about $\mathcal{O}(d\times m)$ parameters per SSM module, where $m$ is the dimension of $k_{t}$, and $d$ is the dimension of $x_{t}$. However, Longhorn demonstrates better performance even with fewer parameteres than Mamba (See Figure 1 (left), Table 3, Table 4). 3. The closed-form solution in Equation 7 does not need any specific initialization. In contrast, Mamba requires careful initialization of the $A$ and $\Delta_{t}$. Comparing to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA). Therefore, it is as fast as existing SSMs. 4. Unlike DeltaNet, which, as noted in Yang et al. (2024), cannot extrapolate beyond the training context, Longhorn demonstrates the capability to successfully extrapolate to contexts up to 16x longer at inference time (See Figure 1 (right)).
Method Online Learning Objective $L_{t}(s)$ (assume $x_{t}\in\mathbb{R}$) Online Update
LA $\left\lVert s-s_{t-1}\right\rVert^{2}-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=s_{t-1}+k_{t}x_{t}$
RetNet $\gamma\left\lVert s-s_{t-1}\right\rVert^{2}+(1-\gamma)\left\lVert s\right\rVert^{2}-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=\gamma s_{t-1}+k_{t}x_{t}$
GLA $(s-s_{t-1})^{\top}\mathrm{diag}(\alpha_{t})(s-s_{t-1})+s^{\top}\mathrm{diag}(1-\alpha_{t})s-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=\mathrm{diag}(\alpha_{t})s_{t-1}+k_{t}x_{t}$
Griffin $\left\lVert\sqrt{\alpha_{t}}\odot(s-s_{t-1})\right\rVert^{2}+\left\lVert\sqrt{1-\alpha_{t}}\odot s\right\rVert^{2}-2\sqrt{1-\alpha_{t}}\odot s\odot i_{t}\odot x_{t}$ $s_{t}=\alpha_{t}\odot s_{t-1}+\sqrt{(1-\alpha_{t})}\odot i_{t}\odot x_{t}$
Longhorn $\left\lVert s-s_{t-1}\right\rVert^{2}+\beta_{t}\left\lVert s^{\top}k_{t}-x_{t}\right\rVert^{2}$ $s_{t}=(1-\Delta_{t}k_{t}^{\odot 2})s_{t-1}+\Delta_{t}k_{t}x_{t}$
$\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$
Table 1: Some of the existing SSMs and their corresponding online learning objectives/updates.
4 Related Work
This section provides a brief summary of recent advances in linear attention (LA) and deep state space models (SSMs).
Linear Attention Models
Several methods have been developed to address the quadratic complexity of the Transformer by making attention linear with respect to context length. In particular, Linformer uses a low-rank approximation of the self-attention by projecting the keys and values into a constant size matrix (instead of scaling with the sequence length)(Wang et al., 2020). Realizing that the main bottleneck causing the quadratic cost in Transformers is the Softmax function, the Linear Transformer replaces the Softmax function with a decomposable similarity function analogous to kernel methods, thereby making the computation linear with respect to sequence length(Katharopoulos et al., 2020). Performer approximates the softmax attention using positive orthogonal random features (Choromanski et al., 2020). More recently, based on the Linear Transformer, the Retentive Network (RetNet) adds additional constant forgetting and rotation (Sun et al., 2023). Gated Linear Attention (GLA) further experiments with learnable forget gates (Yang et al., 2023). Notably, linear attention can be viewed as a fast weight network where a slow net (the model’s parameters) learns to program a changing fast network (e.g., a linear predictor) by adapting its parameters (e.g., the $s_{t}$) online using inputs (Schlag et al., 2021).
State Space Models
Instead of trying to linearize transformers, state space models (SSMs) start with parallelizable linear recurrent networks directly. Initially, the state transition matrix $A$ is assumed to be constant so that the recurrence can be computed in parallel using a convolution (Li et al., 2022; Gu et al., 2021). Subsequent developments include the Diagonal State Space (DSS) model (Gupta et al., 2022), Gated State Space (GSS) models (Mehta et al., 2022), S5 model (Smith et al., 2022), Bidirectional Gated SSM (BiGS)(Wang et al., 2022), H3 model(Fu et al., 2022), and Mamba (Gu & Dao, 2023). In addition, there are also works directly trying to make recurrent networks efficient, which often results in a particular form of SSM as well. This includes Deep Linear Recurrent Units (LRUs)(Orvieto et al., 2023; De et al., 2024), Hierarchically Gated Linear RNN (HGRN)(Qin et al., 2024b; a), and RWKV (Peng et al., 2023; 2024).
Fast Weight Programmer
The concept of a network modifying its own weights in response to inputs is not novel and has historical roots in the Fast-weight Programmer (Schmidhuber, 1992; 1993; Schlag & Schmidhuber, 2017; Schlag et al., 2021). Specifically, these models propose updating a weight matrix $W\in\mathbb{R}^{d\times m}$ using the outer product of two vectors, expressed as $\Delta W=v(x_{t})\otimes k(x_{t})$. This mechanism closely aligns with the principles underlying the Linear Attention model. Our framework builds upon the Fast Weight concept by tailoring the weight update process to address a specific online learning objective, thereby extending its applicability and effectiveness in dynamic learning environments.
Concurrent Work
There are two concurrent works that share similar ideas to ours. Yang et al. (2024) propose a chunk-wise parallel form for scaling the DeltaNet (Schlag et al., 2021) to large-scale language modeling tasks (Yang et al., 2024). The update rule in DeltaNet can be viewed as conducting the gradient step instead of solving for the closed form of the online regression objective, hence the state transition matrix is $A(x_{t})=(I-\beta_{t}k_{t}k_{t}^{\top})$, which is not guaranteed to have eigenvalues $\leq 1$ (the forgetting can even be negative). Therefore, in practice Yang et al. (2024) normalize the key vector $k_{t}$ by its $\ell_{2}$ norm (Schlag et al., 2021), which can be restrictive. In contrast, the closed-form update in Longhorn ensures that the state transition matrix is always stable. In addition, $A(x_{t})$ is a matrix in DeltaNet while we apply a diagonal approximation ($k_{t}^{\odot 2}$). Therefore, Longhorn can leverage either parallel scan (as in Mamba) and chunk-wise parallel training (as in GLA), making it as fast as any existing SSM. In practice, we provide a parallel scan CUDA kernel, so Longhorn serves as a drop-in replacement for Mamba. On the other hand, Sun et al. (2024) propose the Test-Time Training framework, where the update of the state is derived from applying one gradient step of an online regression objective (Sun et al., 2024). To make the model parallelizable, the authors assume that each gradient step at $x_{t}$ ($t\geq 0$) is conducted at state $s_{0}$, hence deriving a parallel form that utilizes matrix multiplication. In contrast, Longhorn solves for the closed-form solution on every token, which potentially leads to a more flexible model.
5 Experiments
We conduct experiments on both synthetic benchmarks and the language modeling task to validate the performance of Longhorn:
  1. 1.
    To test Longhorn’s performance on memorizing associations, we compare it against other SSMs on the synthetic multi-query associative recall benchmark (Arora et al., 2023). We find that Longhorn is the only model that achieves near perfect recall rate when the sequence length scales up to 512 with a hidden dimension of 64).
  2. 2.
    To see how Longhorn performs on language modeling tasks, we first conduct the scaling law experiments using the OpenWebText dataset (Gokaslan & Cohen, 2019), with model sizes of 120M and 350M and context length 1024 or 4096, and then compare the validation perplexity of Longhorn against that of the other SSM models. We find that Longhorn consistently outperforms baseline deep SSM models across context lengths and model sizes.
  3. 3.
    In the end, to see how Longhorn performs on downstream language evaluation tasks, we train a 1.3B language model on the SlimPajama dataset (Soboleva et al., 2023) with 100B tokens, and then compare its downstream evaluation performance across 5 standard benchmarks. We find that Longhorn achieves 1.8x speedup sampling efficiency compared to Mamba, and also outperforms GLA and Mamba on average in downstream evaluation tasks.
5.1 Multi-Query Associative Recall
We first consider the synthetic benchmark Multi-Query Associative Recal (MQAR)l (Arora et al., 2023). The task is the agent will see a sequence of tokens $\{k_{1},v_{1},k_{2},v_{2},\dots,k_{T},v_{T}\}$, where each consecutive two-tokens become a key-value pair. At test time, the agent is provided with multiple $k\sim\{k_{1},\dots k_{T}\}$, the goal is for it to “retrieve” the corresponding values, hence the name multi-query associative recall. Following the original benchmark, we consider the sequence length $T\in\{64,128,256,512\}$ and model dimension (size of the latent embedding of a token) $d\in\{64,128,256,512\}$. We compare against 1) Transformer model (Attention), 2) Based architecture, which combines an SSM with local-attention, where the SSM is derived from the Taylor approximation of the self-attention (Arora et al., 2024), 3) Hyena (Poli et al., 2023), which is a special SSM that adopts long convolution via fast fourier transform, 4) RWKV (Peng et al., 2023), which can be viewed as the division of two SSMs (i.e., $y=a/b$, where $a,b$ are outputs from two SSMs). The state-transition matrix is a scalar, 5) BaseConv (Arora et al., 2023), an SSM that combines linear projection with convolution, and 6) Mamba (Gu & Dao, 2023), the state-of-the-art SSM that has data-dependent $A$ and $B$ (See Equation 3). Each experiment individually searches for the best learning rate from $\{1e-4,4.6e-4,2.2e-3,1e-02\}$ (evenly separated in the log space). Results are summarized in Figure 4.
Figure 4: Comparison of Longhorn against state-of-the-art SSM variants on the MQAR benchmark. The y-axis is the recall rate.
Observation: From the figure, we can see that Longhorn, which is designed to perform the associative recall task by solving the online prediction objective, outperforms existing SSM variants even at the sequence length of 512 and a small model dimension of 64.
5.2 Scaling Law on OpenWebText
In this section we consider language modeling tasks on models with 120M or 350M parameters with 1024 or 4096 context length. We choose the OpenWebText dataset as it is small and serves as an easily accessible benchmark for quick benchmarks. [^2] We adapt the code from the nanoGPT repo, a minimal reproduction of the GPT-2 model. See https://github.com/karpathy/nanoGPT/tree/master. Following the Mamba (Gu & Dao, 2023) experiment setting, the training details are provided in Table 2:
Params n_layers d_model n_heads / d_head Training steps Learning Rate Batch Size Tokens
125M 12 768 12 / 64 4800 6e-4 0.5M tokens 2.5B
350M 24 1024 16 / 64 13500 3e-4 0.5M tokens 7B
Table 2: Training details on OpenWebText.
We consider the following baseline models: LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), Mamba (Gu & Dao, 2023), RWKV (Peng et al., 2023), and GLA (Yang et al., 2023). Then we experiment with 1024 or 4096 context length $T$, which model sizes around 120M or 350M. Results are summarized in Figure 5 and Table 3.
Figure 5: Scaling law with 1024 and 4096 context length on OpenWebText with various SSM models and the LLaMA (strong Transformer) baseline.
Model # Param. (M) Val. Loss ($\downarrow$) # Param. (M) Val. Loss ($\downarrow$)
$T=1024$ $T=4096$ $T=1024$ $T=4096$
RetNet 129.1 3.569 3.492 373.2 3.362 3.227
GLA 123.8 3.381 3.364 361.1 3.018 3.001
RWKV 124.4 3.291 3.276 354.8 2.983 2.931
Mamba 129.2 3.238 3.231 371.5 2.902 2.868
LLaMA 124.4 3.247 3.273 357.7 2.891 2.883
Longhorn 128.6 3.225 3.192 369.8 2.888 2.859
Table 3: Language modeling scaling law against LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), RWKV (Peng et al., 2023), and Mamba (Gu & Dao, 2023). All models are trained on the OpenWebText dataset (Gokaslan & Cohen, 2019). Models vary from around 120M and 350M parameters, with context length 1024 or 4096.
Observation: From the figure and table, we can see that Longhorn consistently outperforms baseline SSMs up to 350M and 4096 context length.
5.3 Large-scale Language Modeling
For the large-scale language modeling task, we followed the experiment setting in GLA (Yang et al., 2023), where a model with around 1.3B parameters is trained on the SlimPajama (Soboleva et al., 2023) dataset over 100B tokens with a batch size of 2M tokens. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with weight decay of 0.01, and apply cosine learning rate decay with a peak learning rate of $43e-$ 4 and the final learning rate of $53e-$ 5. In addition, we use gradient clipping of 1.0. Due to the large size of the experiment, we only compare against LLaMA, Mamba and GLA models, all with a context size of 2048.333Our result is in general better than those reported in the GLA paper, this might due to the fact that we are using a different 100B subset of SlimPajama since the original split of the SlimPajama is not provided.
We evaluate the trained models on a wide range of downstream tasks covering both common-sense reasoning and question-answering. In particular, we consider PIQA (Bisk et al., 2020), HellaSwag (Hella) (Zellers et al., 2019), WinoGrande (Wino) (Sakaguchi et al., 2021), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social Interaction QA (SIQA) (Sap et al., 2019), and Boolean questions (BoolQ) (Clark et al., 2019).
We report the average perplexity across the above eight datasets throughout training in Figure 1 (left). Then we summarize the downstream evaluation results in Table 4.
Model State Size PIQA Hella Wino. ARC-e ARC-c OBQA SIQA BoolQ Avg.
acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$
LLaMA 1024K 55.08 55.36 71.73 59.26 32.19 43.35 45.16 62.13 53.03
GLA 1024K 55.55 49.10 71.12 58.86 28.11 41.67 44.91 59.21 51.07
Mamba 32K 54.21 53.61 71.67 61.05 30.15 43.94 44.18 59.22 52.25
Longhorn 32K 55.78 52.30 71.00 60.63 29.53 43.55 44.68 61.29 52.35
Table 4: Language modeling results against LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2023). All models are trained on the same subset of the SlimPajama dataset with the Mistral tokenizer. The 340M/1.3B models are trained for 15B/100B tokens respectively. State Size is the effective state size of an SSM per layer. For instance, GLA’s state size (1024K) is computed by $md/h$, where the key and value dimensions are $m=1024$ and $d=2048$, and there are 4 heads $h=4$. The individual task performance is via zero-shot. The last column shows the average value over the results on all benchmarks.
Observation: From Figure 1 (left), it is evident that Longhorn not only achieves a lower average perplexity but also improves sampling efficiency by 1.8x compared to Mamba. In other words, Longhorn reaches the same average perplexity with nearly half the training data required by Mamba. From the Table 4, we can see that up to a 1.3B model, Longhorn remains strong among all baseline models and achieves slightly better result than Mamba, even though it has a bit fewer parameters.
5.4 Ablation on Length Extrapolation
Moreover, we evaluate how Longhorn extrapolates to a context length longer than 2048 (training context length) at inference time. In particular, we pick a disjoint validation set from SlimPajama dataset, rearrange it into batches of sequences of length $T\in\{2048,4096,8192,16384,32768\}$, and then evaluate the pretrained model’s perplexity on those sequences. The results are summarized in Figure 1 (right).
Observation: From the figure, we observe that Longhorn successfully extrapolates to contexts up to 16x longer than those used during training, this contrasts with DeltaNet (Yang et al., 2024), which highlights a limitation in that the model cannot extrapolate to longer contexts. Moreover, it is expected that LLaMA, which is a Transformer-based model, fails to extrapolates beyond its training context length.
6 Conclusion and Future Work
In this work, we introduce a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective. We propose a straightforward online regression objective, adopting its implicit closed-form update to define our model, which we refer to as Longhorn. Notably, Longhorn is highly parallelizable and achieves state-of-the-art performance in both synthetic sequence modeling and language modeling tasks. For future research, an intriguing avenue would be exploring other online learning objectives that can be efficiently implemented on modern hardware. Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improve performance. We anticipate similar benefits for Longhorn.

### v4

@@
 Despite extensive research into various design aspects of SSMs, a guiding principle for designing SSMs remains elusive.
-In this work we propose one potential principle.
+In this work, we propose one potential principle.
 We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history into a memory state which is then used by later layers in the network for sequence modeling.
@@
 The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $s_{t}$, which is then passed to later layers for sequence modeling.
-(middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows a certain dynamics that optimizes an online learning objective.
-The recurrent update of $s$ can then be derived by solving the online objective in closed-form or by conducting a gradient step.
+(middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows certain dynamics that optimize an online learning objective.
+The recurrent update of $s$ can then be derived by solving the online objective in closed form or by conducting a gradient step.
 (right) Longhorn’s dynamics is derived from solving an online regression problem, where the goal is to recover $x\in\mathbb{R}$ (if $x\in\mathbb{R}^{d}$ we can run $d$ SSMs in parallel) based on a hint $k\in\mathbb{R}^{m}$ from a state vector $s\in\mathbb{R}^{d}$.
@@
 The hope is that by selecting the right objective, one can save parameters/computation without sacrificing performance.
-In addition, the proposed online learning perspective might lead to better understanding of what SSM layers are doing in a large model.
+In addition, the proposed online learning perspective might lead to a better understanding of what SSM layers are doing in a large model.
 Specifically, we view the recurrent update (a.k.a., the state-transition dynamics) of an SSM model as either a gradient descent step or a closed-form solution to an online learning objective.
@@
 Some variants of SSM models leave the MLP layers unchanged (Sun et al., 2023; Yang et al., 2023; De et al., 2024), while others fuse the SSM layer and the MLP layer into a single unified module (Gu & Dao, 2023).
-But in both cases, the sequence mixing is done by the SSM module and the channel mixing is done by the channel-wise MLP.
-Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of a SSM block for sequence mixing (red), and a MLP block for channel mixing (blue).
+But in both cases, the sequence mixing is done by the SSM module, and the channel mixing is done by the channel-wise MLP.
+Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of an SSM block for sequence mixing (red), and an MLP block for channel mixing (blue).
 SSM: General Form
@@
 A naive linear lay would result in $d\times d\times m$ weights, which is prohibitively large.
-This makes it necessary to impose certain low dimensional structure in $A_{t},B_{t}$, which is what differentiate the existing designs of SSM.
+This makes it necessary to impose certain low-dimensional structures in $A_{t},B_{t}$, which is what differentiates the existing designs of SSM.
 In the following, we provide some examples in the form of Equation 1 for some existing SSM models.
@@
 where $S_{t}\in\mathbb{R}^{d\times m}$, $v(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ are linear mappings of $x_{t}$, and $\otimes$ denote the outer product.
-In practice, one can choose to use $h$ heads as in the multi-head attention to save some computation, where both the $m$ and $d$ dimension are divided into $h$ groups and each group performs its own LA variant.
-The outer product becomes $\mathcal{O}(h*m/h*d/h=md/h)$.
+In practice, one can use $h$ heads as in the multi-head attention to save some computation, where the $m$ and $d$ dimensions are divided into $h$ groups and each group performs its own LA variant.
+The outer product complexity reduces to $\mathcal{O}(h*m/h*d/h=md/h)$.
 But then the effective size of $S_{t}$ also shrinks to $md/h$.
@@
 where $S_{t}\in\mathbb{R}^{d\times m}$ with $m=16$ by default, $\Delta(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ linear mappings of $x_{t}$, and $A\in\mathbb{R}^{d\times m}$ is a data independent (not depending on $x_{t}$ trainable weight matrix.
-We can see that in Mambda both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
+In Mamba, both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
 In practice, Mamba does not use multiple heads as in linear attention variants.
@@
 Example 2.3 (Griffin (De et al., 2024)).
-In Mamba and the linear attention variants, the outer product serves a critical rule of lifting vectors to matrices.
+In Mamba and the linear attention variants, the outer product serves as a critical rule of lifting vectors to matrices.
 The recent Griffin architecture abandons the outer product and performs pure elementwise product:
@@
 where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$.
-This yields smaller memory states, but in practice Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
+This yields smaller memory states, but in practice, Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
 Example 2.4 (RWKV (Peng et al., 2023)).
@@
 where all the vectors are of size $\mathbb{R}^{d}$, and $w>0$ is a trainable weight for controlling the forgetting.
-In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with outer product, which makes it more similar to a LA variant.
+In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with the outer product, which makes it more similar to an LA variant.
 Example 2.5 (HGRN2 (Qin et al., 2024a)).
@@
 3 An Online Learning Perspective For Sequence Mixing
-As demonstrated in the previous section, the design of a state-space model (SSM) depends on the specific selection of $(A_{t},B_{t},*)$, a process that is both intricate and somewhat artisanal.
-In this section, we propose to streamline SSM design through a online learning perspective.
+As demonstrated in the previous section, designing a state-space model (SSM) depends on the specific selection of $(A_{t},B_{t},*)$, which is intricate and somewhat artisanal.
+In this section, we propose to streamline SSM design through an online learning perspective.
 The main idea is to treat the SSM layers as learning modules that learn to compress information along the sequence dimension.
-From this perspective, the SSM layers are learning to learn, such that at the inference time, these layers are still learning (compressing) new information online.
+From this perspective, the SSM layers are learning to learn, such that during the inference time, these layers are still learning (compressing) new information online.
 We begin with an overview of online learning and subsequently demonstrate how SSM can be framed as an online learning problem.
@@
 3.1 SSM as Online Learning
-We advocate to view the recurrence of SSM as solving an online learning problem.
+We advocate viewing the recurrence of SSM as solving an online learning problem.
 In online learning, the agent (the algorithm) picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$.
@@
 Online convex optimization (OCP) yields a principled approach to solving Equation 4 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasiticity” (Mermillod et al., 2013).
-Formally, an online convex programming algorithm update $s_{t}$ by solving a regularized cost function:
+Formally, an online convex programming algorithm updates $s_{t}$ by solving a regularized cost function:
 $s_{t}=\operatorname*{arg\,min}_{s}L_{t}(s),~{}~{}~{}~{}~{}~{}~{}L_{t}(s)=\underbrace{D_{\phi}(s,s_{t-1})}_{\text{stability}}+\underbrace{\beta_{t}\ell_{t}(s)}_{\text{plasticity}},$ (5)
@@
 Here the first term ensures the updated $s$ will be close to the previous $s_{t-1}$, so the agent suffers less from catastrophic forgetting, while the second term ensures the agent is incorporating new knowledge from minimizing the new loss $\ell_{t}(s)$.
-Hence, $\beta_{t}$ controls the trade-off between stability and plasiticity.
+Hence, $\beta_{t}$ controls the trade-off between stability and plasticity.
 3.2 SSM as Online Learning and The Longhorn Architecture
@@
 This provides a unified framework for the existing SSM variants.
-We summarize in Table 1 the online learning interpolation of a number of existing SSM architectures.
-In this work, we explore a highly simplified and natural design called Longhorn guided the online principle (see the last row of Table 1).
-In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is higher dimensional).
-Then, we propose to use the following OCP objective:
-$s_{t}=\operatorname*{arg\,min}_{s}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
-Here, $k_{t}(x_{t})\in\mathbb{R}^{d}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer.
+We summarize in Table 1 the online learning interpretation of several existing SSM architectures.
+In this work, we explore a highly simplified and natural design called Longhorn guided by the online principle (see the last row of Table 1).
+In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is a vector).
+We want to recurrently update hidden states $\{s_{t}\}_{t}$, where $s_{t}\in\mathbb{R}^{m}$ is a vector that aims to summarize the information up to time $t$.
+We posit the following OCP objective for updating $s_{t}$:
+$s_{t}=\operatorname*{arg\,min}_{s\in\mathbb{R}^{m}}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
+Here, $k_{t}(x_{t})\in\mathbb{R}^{m}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer.
 $\beta_{t}(x_{t})$ controls how much new information about $x_{t}$ we want the model to incorporate.
 For instance, $\beta_{t}(x_{t})=0$ implies $s_{t}=s_{t-1}$, while a large $\beta_{t}(x_{t})$ implies the model sacrifices stability for plasticity.
-For simplicity, we will refer to them as $k_{t},\beta_{t}$ from now on.
-From a high-level, Equation 6 is solving an online prediction problem, also known as the associative memory problem (a supervised learning problem), where the goal is to memorize all $(k_{t},x_{t})$ pairs (learn a mapping from $k_{t}$ to $x_{t}$) such that given a key (input) $k$ the model can retrieve (predict) its corresponding value (label) $x$.
-Fortunately, this simple objective gives a closed form solution for $s_{t}$ (a.k.a., the implicit online learning method (Kulis & Bartlett, 2010)).
+For simplicity, we will write $k_{t}=k_{t}(x_{t})$ and $\beta_{t}=\beta_{t}(x_{t})$ from now on.
+From a high-level, Equation 6 is solving an online prediction problem of learning a weight $s$ to predict $x_{t}$ given $k_{t}$ with a linear model $x_{t}\approx s^{\top}k_{t}$.
+It is a supervised formulation of the associative memory problem of memorize $(k_{t},x_{t})$ pairs by learning a mapping from $k_{t}$ to $x_{t}$, such that given a key (input) $k_{t}$ the model can retrieve (predict) its corresponding value (label) $x_{t}$.
+Fortunately, this simple objective gives a closed-form solution for $s_{t}$, which coincides with the implicit online learning method (Kulis & Bartlett, 2010).
 Theorem 3.1.
@@
 ∎
-As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting parallel scan.
-As a result, in practice, we use the diagonal approximation $1-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$.
+As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting a parallel scan.
+As a result, in practice, we use the diagonal approximation $1_{m}-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$ and $1_{m}$ is the $m$-dimensional vector of all ones.
 The final SSM recurrence therefore becomes:
 $s_{t}=\underbrace{(1-\Delta_{t}k_{t}^{\odot 2})}_{A_{t}(x_{t})}\odot s_{t-1}+\underbrace{\Delta_{t}k_{t}x_{t}}_{B_{t}(x_{t})}.$ (8)
-When $x_{t}\in\mathbb{R}^{d}$ is a higher dimensional vector, following the Mamba design, we use $d$ parallel SSMs.
+Following the Mamba design, when $x_{t}\in\mathbb{R}^{d}$ is a vector, we use $d$ parallel SSMs.
 As a result, $x_{t},\beta_{t},\Delta_{t}\in\mathbb{R}^{d},k_{t}\in\mathbb{R}^{m},S_{t}\in\mathbb{R}^{d\times m}$, Longhorn’s update becomes:
@@
 We also provide an efficient CUDA kernel for it.
-The full inference time algorithm is provided in Algorithm 1.
+The full inference-time algorithm is provided in Algorithm 1.
 One can compare Equation 9 to Equation 3 and other SSM in Section 2.
@@
 This saves about $\mathcal{O}(d\times m)$ parameters per SSM module, where $m$ is the dimension of $k_{t}$, and $d$ is the dimension of $x_{t}$.
-However, Longhorn demonstrates better performance even with fewer parameteres than Mamba (See Figure 1 (left), Table 3, Table 4).
+However, Longhorn demonstrates better performance even with fewer parameters than Mamba (See Figure 1 (left), Table 3, Table 4).
 3.
@@
 In contrast, Mamba requires careful initialization of the $A$ and $\Delta_{t}$.
-Comparing to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA).
+Compared to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA).
 Therefore, it is as fast as existing SSMs.
@@
 For future research, an intriguing avenue would be exploring other online learning objectives that can be efficiently implemented on modern hardware.
-Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improve performance.
+Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improves performance.
 We anticipate similar benefits for Longhorn.

### v5

@@
 Abstract
-The most fundamental capability of modern AI methods such as Large Language Models (LLMs) is the ability to predict the next token in a long sequence of tokens, known as “sequence modeling.” Although the Transformers model is the current dominant approach to sequence modeling, its quadratic computational cost with respect to sequence length is a significant drawback.
-State-space models (SSMs) offer a promising alternative due to their linear decoding efficiency and high parallelizability during training.
-However, existing SSMs often rely on seemingly ad hoc linear recurrence designs.
+Modern large language models are built on sequence modeling via next-token prediction.
+While the Transformer remains the dominant architecture for sequence modeling, its quadratic decoding complexity in sequence length poses a major limitation.
+State-space models (SSMs) present a competitive alternative, offering linear decoding efficiency while maintaining parallelism during training.
+However, most existing SSMs rely on linear recurrence designs that appear somewhat ad hoc.
 In this work, we explore SSM design through the lens of online learning, conceptualizing SSMs as meta-modules for specific online learning problems.
-This approach links SSM design to formulating precise online learning objectives, with state transition rules derived from optimizing these objectives.
-Based on this insight, we introduce a novel deep SSM architecture based on the implicit update for optimizing an online regression objective.
-Our experimental results show that our models outperform state-of-the-art SSMs, including the Mamba model, on standard sequence modeling benchmarks and language modeling tasks.
-[^1] Code is available at https://github.com/Cranial-XIX/longhorn.git.
-Figure 1: (left) The average perplexity across eight downstream datasets for GLA, Mamba and Longhorn (1.3B model) over seen tokens on SlimPajama.
+This approach links SSM design to formulating precise online learning objectives, with state transition rules derived from solving these objectives.
+Based on this insight, we introduce a novel deep SSM architecture, Longhorn, whose update resembles the closed-form solution for solving the online associative recall problem.
+Our experimental results show that Longhorn outperforms state-of-the-art SSMs, including the Mamba model, on standard sequence modeling benchmarks, language modeling, and vision tasks.
+Specifically, Longhorn achieves a 1.8x improvement in sample efficiency compared to Mamba, and can extrapolate over contexts that are up to 16x longer during inference.
+Figure 1: (left) The average perplexity on eight downstream datasets for GLA, Mamba, and Longhorn (1.3B model) over seen tokens on SlimPajama.
 Longhorn leads to a 1.8x speed up in sampling efficiency.
-(right) Longhorn, pretrained with 2048 context length, successfully extrapolates to longer context (up to 16x of the trained context).
+(right) Longhorn, pretrained with 2048 context length, extrapolates up to 16x longer context at inference.
 1 Introduction
@@
 However, its utility is constrained by the quadratic growth in training and decoding costs with increasing sequence length.
-Despite various optimizations such as efficient decoding (Chen et al., 2023; Kuperman & Dyke, 2011), KV-cache compression (DeepSeek-AI & Dai, 2024), and memory efficient implementation (Dao et al., 2022), scaling Transformers for autonomous and continual use with an infinite (or very long) context window remains a challenge.
-Recent advancements in linear attention models (Katharopoulos et al., 2020) and state-space models (SSMs)(Gu et al., 2021) have demonstrated their potential.
+Despite various optimizations such as efficient decoding (e.g., Chen et al., 2023; Kuperman & Dyke, 2011), KV-cache compression (e.g., DeepSeek-AI & Dai, 2024), and memory efficient implementation (e.g., Dao et al., 2022), it remains challenging to scale Transformers for autonomous and continual use with an infinite (or very long) context window.
+Recent advances in linear attention models (Katharopoulos et al., 2020) and state-space models (SSMs) (Gu et al., 2021) have demonstrated their potential.
 These models are specialized recurrent neural networks capable of efficiently computing outputs in parallel when input tokens are provided simultaneously during training, thus avoiding the inefficiencies of traditional backpropagation through time.
@@
 Initially, these models underperformed compared to Transformers.
-However, recent SSMs (Gu & Dao, 2023; Yang et al., 2023; Peng et al., 2024; De et al., 2024; Beck et al., 2024) have achieved performance parity with Transformers in language modeling tasks.
+However, recent SSMs (e.g., Gu & Dao, 2023; Yang et al., 2023; Peng et al., 2024; De et al., 2024; Beck et al., 2024) have achieved performance parity with Transformers in language modeling tasks.
 Despite extensive research into various design aspects of SSMs, a guiding principle for designing SSMs remains elusive.
 In this work, we propose one potential principle.
-We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history into a memory state which is then used by later layers in the network for sequence modeling.
+We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history online into a memory state which is then used by later layers in the network for sequence modeling.
 From this perspective:
@@
 The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $S_{t}$, which is then passed to later layers for sequence modeling.
-(middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows certain dynamics that optimize an online learning objective.
-The recurrent update of $s$ can then be derived by solving the online objective in closed form or by conducting a gradient step.
-(right) Longhorn’s dynamics is derived from solving an online regression problem, where the goal is to recover $x\in\mathbb{R}$ (if $x\in\mathbb{R}^{d}$ we can run $d$ SSMs in parallel) based on a hint $k\in\mathbb{R}^{m}$ from a state vector $s\in\mathbb{R}^{d}$.
-The closed-form solution to the online regression problem corresponds to the implicit online learning update, where $\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$.
-Please see the details in Section 3.
+(middle) Sequence mixing can be seen as an online learning problem, where the SSM state $S_{t}$ optimizes an online objective.
+The recurrent update of $S_{t}$ is derived either by solving this objective in closed form or via a proximal update.
+(right) Longhorn’s update solves online associative recall, where the goal is to recover $x\in\mathbb{R}^{d}$ based on a hint $k\in\mathbb{R}^{m}$ from a state matrix $S\in\mathbb{R}^{d\times m}$.
+Longhorn’s update corresponds to the implicit online learning solution, where $A_{t}=1_{d\times m}-\varepsilon_{t}\otimes k^{\otimes 2}$ and $B_{t}=(\varepsilon\odot x_{t})\otimes k_{t}$, and $\varepsilon_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$.
+See the details in Section 3 and Algorithm 1.
 As a result, we can draw inspiration from online learning and confine the design choices of SSMs to reflect those learning dynamics that solve specific online prediction problems.
-The hope is that by selecting the right objective, one can save parameters/computation without sacrificing performance.
-In addition, the proposed online learning perspective might lead to a better understanding of what SSM layers are doing in a large model.
-Specifically, we view the recurrent update (a.k.a., the state-transition dynamics) of an SSM model as either a gradient descent step or a closed-form solution to an online learning objective.
-We provide the objectives for some existing SSMs in Table 1.
+The aim is that, by optimizing for the right objective, the model can achieve superior performance with fewer parameters or reduced computational costs.
+Furthermore, this online learning perspective may offer deeper insights into the function of SSM layers in large models.
+In particular, the recurrent update (i.e., state-transition dynamics) of an SSM can be interpreted as a proximal update step or a closed-form solution to an online learning objective.
+We outline the corresponding objectives for several existing SSMs in Table 4.
+One significant advantage of viewing SSMs through the lens of online learning is their ability to adapt post-training during deployment, allowing them to process arbitrarily long data sequences at inference time.
 Based on this insight, we propose a simple yet effective architecture (Longhorn), derived from the implicit closed-form update of an online associative recall problem.
-The closed-form update naturally leads to a stable recurrent form without a manually designed gating mechanism.
-So Longhorn does not need a separately parameterized forget gate, which saves parameters when the state size is large.
-We demonstrate that Longhorn performs comparably to or better than state-of-the-art SSMs like Mamba (Gu & Dao, 2023) on both synthetic and large-scale sequence modeling tasks.
+The closed-form update naturally leads to a stable recurrent form without a manually designed gating mechanism, automatically balancing forgetting and learning.
+Thus Longhorn does not need a separately parameterized forget gate, which saves parameters when the state size is large.
+We demonstrate that Longhorn performs comparably to or better than state-of-the-art SSMs like Mamba (Gu & Dao, 2023) on synthetic and large-scale sequence modeling tasks.
 In particular, Longhorn outperforms Mamba at the size of 1.3B-parameter when trained on 100B tokens from the SlimPajama dataset (Soboleva et al., 2023).
 To summarize, our contributions are:
-1) Theoretical Framework: We propose a novel framework that views SSM’s recurrent update as solving online learning objectives.
+1) Theoretical Framework: We propose a novel framework that views SSMs’ recurrent update as solving online learning objectives.
 As a result, the design of SSMs reduces to the design of the online learning objectives.
@@
 2) Empirical Results: Longhorn demonstrates better performance than existing SSMs including Mamba, across both synthetic associative recall tasks and the large-scale language modeling task.
-Moreover, it achieves 1.8x enhancement in sampling efficiency compared to Mamba (See Figure 1 (left)).
+Moreover, it achieves 1.8x improvement in sample efficiency compared to Mamba (See Figure 1 (left)).
 Longhorn’s training speed is as fast as Mamba, as we only replace the SSM module in the Mamba architecture with Longhorn’s recurrence.
@@
 4:     1) Receive input $x_{t}\in\mathbb{R}^{d}$.
-5:     2) Compute the query $q_{t}$, key $k_{t}$ and $\beta_{t}$ (as in objective $||s-s_{t-1}||^{2}+\beta_{t}||s^{\top}k_{t}-x_{t}||^{2}$):
+5:     2) Compute the query $q_{t}$, key $k_{t}$ and $\beta_{t}$:
 $q_{t}=W_{q}x_{t}\in\mathbb{R}^{m},$ $k_{t}=W_{k}x_{t}\in\mathbb{R}^{m},$ $\beta_{t}=\text{Sigmoid}(W_{\beta}x_{t})\in(0,1)^{d}.$
 6:     3) Update the memory state $S_{t}\in\mathbb{R}^{d\times m}$ via
-$S_{t}=\big{(}1-\Delta_{t}\otimes k_{t}^{\odot 2}\big{)}\odot S_{t-1}+\big{(}\Delta_{t}\odot x_{t}\big{)}\otimes k_{t},$
-where $\Delta_{t}$ is the step size:
-$\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})\in(0,1)^{d}.$
+$S_{t}=\big{(}1_{d\times m}-\varepsilon_{t}\otimes k_{t}^{\odot 2}\big{)}\odot S_{t-1}+\big{(}\varepsilon_{t}\odot x_{t}\big{)}\otimes k_{t},$ $\varepsilon_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})\in(0,1)^{d}.$
 7:     4) Compute the output $o_{t}=S_{t}q_{t}\in\mathbb{R}^{d}$.
@@
 But in both cases, the sequence mixing is done by the SSM module, and the channel mixing is done by the channel-wise MLP.
-Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of an SSM block for sequence mixing (red), and an MLP block for channel mixing (blue).
+Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi_{i}(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of an SSM block for sequence mixing (red), and an MLP block for channel mixing (blue).
 SSM: General Form
@@
 $S_{t}=A(x_{t})*S_{t-1}+B(x_{t}),$ $\forall t\in\{1,\ldots,T\},$ $S_{0}=0,$ (1)
-where $x_{t}$ is the input at time $t$, $S_{t}$ is the model’s state, $A_{t},B_{t}\colon\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are some functions and * is a multiplication operation of choice, such as Hadamard product or matrix product.
+where $x_{t}$ is the input at time $t$, $S_{t}$ is the model’s state, $A_{t},B_{t}\colon\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are some functions of the input and * is a multiplication operation of choice, such as Hadamard product or matrix product.
 Given the state $S_{t}$, SSMs often give the output token at the next layer via a gated linear unit (GLU) (Dauphin et al., 2017):
@@
 where we first get $o_{t}$ via a state-dependent linear projection on $S_{t}$, which is then fed into a subsequent channel mixing gated linear unit (blue in Figure 3), where $\sigma(\cdot)$ is a non-linear activation function.
+A key feature of this design in Equation 1 is that $S_{t}$ has a linear recurrence, i.e., $S_{t}$ is linear in $S_{t-1}$.
 Figure 3: Mamba Block
-A key feature of this design in Equation 1 is that the update of $S_{t}$ is a linear recurrence relation.
-That is, $S_{t}$ is a linear function of $S_{t-1}$.
 Crucially, this allows us to express all $S_{t}$ in an explicit form that can be calculated in parallel: when all $\bm{x}=\{x_{t}\}_{t}$ are available as in the training phase, $\{S_{t}\}_{t}$ can be written into
@@
 Here $\prod$ denotes the product induced by multiplication operator *.
-The resulting cumulative product $\overline{A}_{t^{\prime}\to t}$ can be implemented efficiently in parallel with the prefix scan algorithm (Harris et al., 2007), which only requires a sublinear complexity in terms of sequence length (e.g., $\mathcal{O}(\log t)$).
+The resulting cumulative product $\overline{A}_{t^{\prime}\to t}$ can be implemented efficiently in parallel with the prefix scan algorithm (e.g., Harris et al., 2007), which only requires $\mathcal{O}(\log T)$ ($T$ is the sequence length) parallel operations.
 From now on, we will abbreviate $A(x_{t})$ and $B(x_{t})$ as $A_{t}$ and $B_{t}$, respectively.
 Designs of ($A_{t}$, $B_{t}$, *)
-Existing variants of SSMs mainly differs in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence.
-A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed be $m$ times of the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence.
+Existing variants of SSMs mainly differ in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence.
+A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed to be $m$ times the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence.
 This makes the architecture design of $A_{t},B_{t}$, both mapping $\mathbb{R}^{d}$ to $\mathbb{R}^{d\times m}$ challenging.
-A naive linear lay would result in $d\times d\times m$ weights, which is prohibitively large.
-This makes it necessary to impose certain low-dimensional structures in $A_{t},B_{t}$, which is what differentiates the existing designs of SSM.
-In the following, we provide some examples in the form of Equation 1 for some existing SSM models.
-Example 2.1 (Linear Attention Variants).
-Linear Attention (LA) (Katharopoulos et al., 2020), Retention Network (RetNet) (Sun et al., 2023), and Gated Linear Attention (GLA) (Yang et al., 2023) all assume $A_{t},B_{t}$ yield rank-1 (or even constant) outputs:
-$S_{t}=A_{t}\odot S_{t-1}+v(x_{t})\otimes k(x_{t}),~{}~{}~{}~{}\text{with}~{}~{}\begin{cases}A_{t}=1&\quad\text{(LA)}\\ A_{t}=c\in[0,1]&\quad\text{(RetNet)}\\ A_{t}=1\otimes\alpha(x_{t})&\quad\text{(GLA)}\end{cases},$
-where $S_{t}\in\mathbb{R}^{d\times m}$, $v(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ are linear mappings of $x_{t}$, and $\otimes$ denote the outer product.
-In practice, one can use $h$ heads as in the multi-head attention to save some computation, where the $m$ and $d$ dimensions are divided into $h$ groups and each group performs its own LA variant.
-The outer product complexity reduces to $\mathcal{O}(h*m/h*d/h=md/h)$.
-But then the effective size of $S_{t}$ also shrinks to $md/h$.
-Example 2.2 (Mamba (Gu & Dao, 2023)).
-The Mamba architecture is derived by discretizing a continuous linear dynamics. Its discretized update is:
-$\begin{split}S_{t}&=A_{t}\odot S_{t-1}+B_{t},~{}~{}~{}~{}\text{where}\\ A_{t}&=\exp(A\odot(\Delta(x_{t})\otimes 1)),\qquad B_{t}=(\Delta(x_{t})\odot x_{t})\otimes k(x_{t}).\end{split}$ (3)
-where $S_{t}\in\mathbb{R}^{d\times m}$ with $m=16$ by default, $\Delta(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ linear mappings of $x_{t}$, and $A\in\mathbb{R}^{d\times m}$ is a data independent (not depending on $x_{t}$ trainable weight matrix.
-In Mamba, both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
-In practice, Mamba does not use multiple heads as in linear attention variants.
-Perhaps the main reason is that given a fixed $m$ and $d$, the largest memory state will be with $h=1$ (as the effective size of $S_{t}$ is $md/h$).
-In addition, Mamba’s output is $o_{t}=C(x_{t})S_{t}+D_{t}\odot x_{t}$, which has an additional residual part $D_{t}\odot x_{t}$.
-Example 2.3 (Griffin (De et al., 2024)).
-In Mamba and the linear attention variants, the outer product serves as a critical rule of lifting vectors to matrices.
-The recent Griffin architecture abandons the outer product and performs pure elementwise product:
-$s_{t}=a(x_{t})\odot s_{t-1}+\sqrt{1-a(x_{t})}\odot i(x_{t})\odot x_{t},$
-where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$.
-This yields smaller memory states, but in practice, Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
-Example 2.4 (RWKV (Peng et al., 2023)).
-The original RWKV also performs elementwise recurrence. It maintains a state of ratio form $s_{t}=u_{t}/z_{t}$, where $u_{t},z_{t}$ are updated separately by two SSMs:
-$s_{t}$ $=u_{t}/z_{t}$
-$u_{t}$ $=\exp(-w)\cdot u_{t-1}+\exp(k(x_{t}))\odot v(x_{t}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}z_{t}=\exp(-w)\cdot z_{t-1}+\exp(k(x_{t})),$
-where all the vectors are of size $\mathbb{R}^{d}$, and $w>0$ is a trainable weight for controlling the forgetting.
-In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with the outer product, which makes it more similar to an LA variant.
-Example 2.5 (HGRN2 (Qin et al., 2024a)).
-The Gated Linear RNNs with State Expansion (HGRN2) model is represented with the following recurrence:
-$S_{t}=(1\otimes f(x_{t}))\odot S_{t-1}+i(x_{t})\otimes(1-f(x_{t})).$
-Here, $f(x_{t})\in[0,1]$ is the forget gate, $(1-f(x_{t}))$ is the input gate, and $i(x_{t})$ is the input vector. HGRN2 thus resembles an RNN.
+A naive linear mapping would result in $d\times d\times m$ weights, which is prohibitively large.
+This makes it necessary to impose certain low-dimensional structures in $A_{t},B_{t}$, which is the main difference from existing designs of SSMs.
+In Appendix A, we summarize some existing deep SSM models in the form of Equation 1.
 3 An Online Learning Perspective For Sequence Mixing
@@
 We advocate viewing the recurrence of SSM as solving an online learning problem.
-In online learning, the agent (the algorithm) picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$.
+In online learning, the agent picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$.
 The goal is to minimize
 $\min_{\{s_{t}\}}\sum_{t}\ell_{t}(s_{t}).$ (3)
-For instance, consider online linear prediction, where at each step the agent is given an input-label pair $(x_{t},y_{t})$ and $\ell_{t}(s_{t})=\frac{1}{2}||s_{t}^{\top}x_{t}-y_{t}||^{2}$ is the $\ell_{2}$ regression loss, then the problem becomes an online regression problem, and the goal is to successfully predict $y_{t}$ given $x_{t}$ at future time steps.
-Online convex optimization (OCP) yields a principled approach to solving Equation 4 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasiticity” (Mermillod et al., 2013).
+For instance, consider online linear prediction, where at each step the agent is given an input-label pair $(x_{t},y_{t})$ and $\ell_{t}(s_{t})=\frac{1}{2}||s_{t}^{\top}x_{t}-y_{t}||^{2}$ is the $\ell_{2}$ regression loss, then the problem becomes an online regression problem, and the goal is to successfully predict $y_{t}$ given $x_{t}$ at future time steps, with the key feature that the prediction ($s_{t}$) can change with each new data point.
+Online convex programming (OCP) (e.g., Zinkevich, 2003) yields a principled approach to solving Equation 3 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasticity” (e.g., Mermillod et al., 2013).
 Formally, an online convex programming algorithm updates $s_{t}$ by solving a regularized cost function:
@@
 Hence, $\beta_{t}$ controls the trade-off between stability and plasticity.
-3.2 SSM as Online Learning and The Longhorn Architecture
-Under the online learning framework, the design of SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 5.
+3.2 The Longhorn Architecture
+Under the online learning framework, the design of an SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 4.
 This provides a unified framework for the existing SSM variants.
-We summarize in Table 1 the online learning interpretation of several existing SSM architectures.
+We summarize in Table 4 in Appendix B the online learning interpretation of several existing SSM architectures.
 In this work, we explore a highly simplified and natural design called Longhorn guided by the online principle (see the last row of Table 4).
-In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is a vector).
-We want to recurrently update hidden states $\{s_{t}\}_{t}$, where $s_{t}\in\mathbb{R}^{m}$ is a vector that aims to summarize the information up to time $t$.
+In particular, we consider $\{(k_{t},x_{t})\}_{t}$ as the input stream, where $k_{t}\in\mathbb{R}^{m}$ and $x_{t}\in\mathbb{R}^{d}$ are the key-value pairs, just as in the Transformer model (Vaswani et al., 2017).
+In practice, as in Mamba (Gu & Dao, 2023), $k_{t}=W_{k}x_{t}\in\mathbb{R}^{m}$, where $W_{k}\in\mathbb{R}^{m\times d}$, is a linear mapping from $x_{t}$.
+We want to recurrently update hidden states $\{S_{t}\}_{t}$, where $S_{t}\in\mathbb{R}^{d\times m}$ is a matrix that summarizes the information up to time $t$.
 We posit the following OCP objective for updating $S_{t}$:
-$s_{t}=\operatorname*{arg\,min}_{s\in\mathbb{R}^{m}}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
-Here, $k_{t}(x_{t})\in\mathbb{R}^{m}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer.
-$\beta_{t}(x_{t})$ controls how much new information about $x_{t}$ we want the model to incorporate.
-For instance, $\beta_{t}(x_{t})=0$ implies $s_{t}=s_{t-1}$, while a large $\beta_{t}(x_{t})$ implies the model sacrifices stability for plasticity.
-For simplicity, we will write $k_{t}=k_{t}(x_{t})$ and $\beta_{t}=\beta_{t}(x_{t})$ from now on.
-From a high-level, Equation 6 is solving an online prediction problem of learning a weight $s$ to predict $x_{t}$ given $k_{t}$ with a linear model $x_{t}\approx s^{\top}k_{t}$.
+$S_{t}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d\times m}}\left\{||S-S_{t-1}||^{2}_{\text{F}}+||Sk_{t}-x_{t}||^{2}_{\mathrm{diag}(\beta_{t})}\right\}.$ (5)
+Here, $||\cdot||_{\text{F}}$ denotes the Frobenius norm of a matrix, $\beta_{t}\in\mathbb{R}^{d}$ is a vector controlling how much new information about $x_{t}$ we want the model to incorporate for $S_{t}$.
+For instance, $\beta_{t,i}=0$ implies $S_{t,i}=S_{t-1,i}$ (i.e., the $i$-th row of $S$ remains unchanged), while a large $\beta_{t,i}$ implies the model empties some part of $S_{i}$ for incorporating $x_{t,i}$.
+From a high-level perspective, Equation 5 is solving an online prediction problem of learning a weight matrix $S$ to predict $x_{t}$ given $k_{t}$ with a linear model $x_{t}\approx S^{\top}k_{t}$.
 It is a supervised formulation of the associative memory problem of memorize $(k_{t},x_{t})$ pairs by learning a mapping from $k_{t}$ to $x_{t}$, such that given a key (input) $k_{t}$ the model can retrieve (predict) its corresponding value (label) $x_{t}$.
-Fortunately, this simple objective gives a closed-form solution for $s_{t}$, which coincides with the implicit online learning method (Kulis & Bartlett, 2010).
+The objective in Equation 5 is motivated by the observation that the self-attention layer of the Transformer exhibits a form of online associative recall (often referred to as the induction head property) (Olsson et al., 2022).
+This capability has been shown to underpin the model’s ability to perform in-context learning (Brown, 2020).
+To explain the connection, in-context learning refers to the model’s ability, during inference, to generalize from a set of provided ${(k,x)}$ (question-answer) pairs and apply this understanding to a new question.
+This closely parallels associative recall, where the model retrieves relevant information from past interactions to address new inputs.
+Fortunately, this simple objective gives a closed-form solution for $S_{t}$, which coincides with the implicit online learning method (e.g., Kulis & Bartlett, 2010), according to Theorem 6 (We provide the proof in Appendix C):
 Theorem 3.1.
 The closed form solution for $S_{t}$ for objective in Equation 5 is
-$s_{t}=(I-\Delta_{t}k_{t}k_{t}^{\top})s_{t-1}+\Delta_{t}k_{t}x_{t},~{}~{}~{}\text{where}~{}\Delta_{t}=\frac{\beta_{t}}{1+\beta_{t}k_{t}^{\top}k_{t}}\in[0,\infty).$ (7)
-Proof.
-As the objective in Equation 6 is in a quadratic form with respect to $s$, there is a unique minimum:
-$\nabla L_{t}(s)=0$ $\Longleftrightarrow(s-s_{t-1})+\beta_{t}(s^{\top}k_{t}-x_{t})k_{t}=0$
-$\Longleftrightarrow(I+\beta_{t}k_{t}k_{t}^{\top})s=s_{t-1}+\beta_{t}k_{t}x_{t}$
-$\underbrace{\Longleftrightarrow}_{(3)}s=\bigg{(}I-\frac{\beta_{t}}{1+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}\beta_{t}k_{t}x_{t}$
-$\Longleftrightarrow\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\frac{(I+\beta_{t}k_{t}^{\top}k_{t}-\beta_{t}k_{t}k_{t}^{\top})\beta_{t}k_{t}x_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}$
-$\underbrace{\Longleftrightarrow}_{(5)}\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\frac{\beta_{t}k_{t}x_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}$
-(3) is derived from the fact that $(I+\beta_{t}k_{t}k_{t}^{\top})^{-1}=(I-\frac{\beta_{t}k_{t}k_{t}^{\top}}{I+\beta k_{t}^{\top}k_{t}})$ by the Sherman–Morrison formula.
-(5) is derived by noticing that $k_{t}^{\top}k_{t}k_{t}x_{t}-k_{t}k_{t}^{\top}k_{t}x_{t}=0$.
-∎
+$S_{t,i}=(I-\varepsilon_{t,i}k_{t}k_{t}^{\top})S_{t-1,i}+\varepsilon_{t,i}k_{t}x_{t,i},~{}~{}~{}\text{where}~{}\varepsilon_{t,i}=\frac{\beta_{t,i}}{1+\beta_{t,i}k_{t}^{\top}k_{t}}\in[0,\infty).$ (6)
+Here, $S_{t,i}$ refers to the $i$-th row of $S_{t}$, $\beta_{t,i}$ refers to the $i$-th element of $\beta_{t}$.
 As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting a parallel scan.
-As a result, in practice, we use the diagonal approximation $1_{m}-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$ and $1_{m}$ is the $m$-dimensional vector of all ones.
-The final SSM recurrence therefore becomes:
-$s_{t}=\underbrace{(1-\Delta_{t}k_{t}^{\odot 2})}_{A_{t}(x_{t})}\odot s_{t-1}+\underbrace{\Delta_{t}k_{t}x_{t}}_{B_{t}(x_{t})}.$ (8)
-Following the Mamba design, when $x_{t}\in\mathbb{R}^{d}$ is a vector, we use $d$ parallel SSMs.
-As a result, $x_{t},\beta_{t},\Delta_{t}\in\mathbb{R}^{d},k_{t}\in\mathbb{R}^{m},S_{t}\in\mathbb{R}^{d\times m}$, Longhorn’s update becomes:
-$S_{t}=A_{t}\odot S_{t-1}+B_{t},~{}~{}\text{where}~{}~{}A_{t}=(1-\Delta_{t}\otimes k_{t}^{\odot 2}),~{}~{}B_{t}=(\Delta_{t}\odot x_{t})\otimes k_{t}.$ (9)
-The final architecture of Longhorn follows Mamba strictly (Figure 3), and we only replace the SSM block with the Longhorn’s recurrence.
+As a result, in practice, we use the diagonal approximation $1_{m}-\varepsilon_{t,i}k_{t}^{\odot 2}$ in place of $I-\varepsilon_{t,i}k_{t}k_{t}^{\top}$, where $a^{\odot 2}=a\odot a$ and $1_{m}$ is the $m$-dimensional all-one vector.
+Following Mamba (Gu & Dao, 2023) and Transformer (Vaswani et al., 2017), we make $k_{t}=W_{k}x_{t}\in\mathbb{R}^{m}$ and $\beta_{t}=\sigma(W_{\beta}x_{t})\in\mathbb{R}^{d}$ (both are functions of $x_{t}$), where the activation $\sigma$ (the Sigmoid function) is to ensure that $\beta_{t}$ is positive and bounded.
+In summary, the final Longhorn update of $S_{t}$ becomes:
+$S_{t}=A_{t}\odot S_{t-1}+B_{t},~{}~{}\text{where}~{}~{}A_{t}=(1_{d\times m}-\varepsilon_{t}\otimes k_{t}^{\odot 2}),~{}~{}B_{t}=(\varepsilon_{t}\odot x_{t})\otimes k_{t}.$ (7)
+The final architecture of Longhorn follows Mamba strictly (Figure 3), except that we replace the SSM block with Longhorn’s recurrence.
 We also provide an efficient CUDA kernel for it.
 The full inference-time algorithm is provided in Algorithm 1.
-One can compare Equation 9 to Equation 3 and other SSM in Section 2.
-Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forget gate is naturally derived from the key vector, i.e., $(1-k^{\odot 2}\otimes\Delta_{t})$.
-Advantages of Longhorn: 1.
+One can compare Equation 7 to Equation 8 and other SSMs in Appendix A.
+Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forgetting gate is naturally derived from the key vector, i.e., $(1_{d\times m}-\varepsilon_{t}\otimes k_{t}^{\odot 2})$.
+Advantages of Longhorn
+  1. 1.
     While we can derive the learning objective for some of the existing SSMs, Longhorn is the first SSM designed for explicitly solving an online regression problem.
-2.
+  2. 2.
     Longhorn does not require a specific forget gate (e.g., $\alpha_{t}$ in GLA or $A$ matrix in Mamba).
@@
 However, Longhorn demonstrates better performance even with fewer parameters than Mamba (See Figure 1 (left), Table 1, Table 2).
-3.
-The closed-form solution in Equation 7 does not need any specific initialization.
-In contrast, Mamba requires careful initialization of the $A$ and $\Delta_{t}$.
-Compared to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA).
-Therefore, it is as fast as existing SSMs.
-4.
-Unlike DeltaNet, which, as noted in Yang et al.
-(2024), cannot extrapolate beyond the training context, Longhorn demonstrates the capability to successfully extrapolate to contexts up to 16x longer at inference time (See Figure 1 (right)).
-Method Online Learning Objective $L_{t}(s)$ (assume $x_{t}\in\mathbb{R}$) Online Update
-LA $\left\lVert s-s_{t-1}\right\rVert^{2}-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=s_{t-1}+k_{t}x_{t}$
-RetNet $\gamma\left\lVert s-s_{t-1}\right\rVert^{2}+(1-\gamma)\left\lVert s\right\rVert^{2}-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=\gamma s_{t-1}+k_{t}x_{t}$
-GLA $(s-s_{t-1})^{\top}\mathrm{diag}(\alpha_{t})(s-s_{t-1})+s^{\top}\mathrm{diag}(1-\alpha_{t})s-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=\mathrm{diag}(\alpha_{t})s_{t-1}+k_{t}x_{t}$
-Griffin $\left\lVert\sqrt{\alpha_{t}}\odot(s-s_{t-1})\right\rVert^{2}+\left\lVert\sqrt{1-\alpha_{t}}\odot s\right\rVert^{2}-2\sqrt{1-\alpha_{t}}\odot s\odot i_{t}\odot x_{t}$ $s_{t}=\alpha_{t}\odot s_{t-1}+\sqrt{(1-\alpha_{t})}\odot i_{t}\odot x_{t}$
-Longhorn $\left\lVert s-s_{t-1}\right\rVert^{2}+\beta_{t}\left\lVert s^{\top}k_{t}-x_{t}\right\rVert^{2}$ $s_{t}=(1-\Delta_{t}k_{t}^{\odot 2})s_{t-1}+\Delta_{t}k_{t}x_{t}$
-$\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$
-Table 1: Some of the existing SSMs and their corresponding online learning objectives/updates.
+  3. 3.
+    The closed-form solution in Equation 6 does not need any specific initialization. In contrast, Mamba requires a special careful initialization of the $A$ and $\varepsilon_{t}$.
+  4. 4.
+    Unlike DeltaNet (Yang et al., 2024), which struggles to extrapolate beyond training contexts, Longhorn successfully extrapolates to contexts 16x longer than it was trained for (Figure 1 (right)).
 4 Related Work
-This section provides a brief summary of recent advances in linear attention (LA) and deep state space models (SSMs).
+This section provides a summary of recent advances in linear attention and state space models.
 Linear Attention Models
-Several methods have been developed to address the quadratic complexity of the Transformer by making attention linear with respect to context length.
-In particular, Linformer uses a low-rank approximation of the self-attention by projecting the keys and values into a constant size matrix (instead of scaling with the sequence length)(Wang et al., 2020).
-Realizing that the main bottleneck causing the quadratic cost in Transformers is the Softmax function, the Linear Transformer replaces the Softmax function with a decomposable similarity function analogous to kernel methods, thereby making the computation linear with respect to sequence length(Katharopoulos et al., 2020).
-Performer approximates the softmax attention using positive orthogonal random features (Choromanski et al., 2020).
-More recently, based on the Linear Transformer, the Retentive Network (RetNet) adds additional constant forgetting and rotation (Sun et al., 2023).
-Gated Linear Attention (GLA) further experiments with learnable forget gates (Yang et al., 2023).
-Notably, linear attention can be viewed as a fast weight network where a slow net (the model’s parameters) learns to program a changing fast network (e.g., a linear predictor) by adapting its parameters (e.g., the $s_{t}$) online using inputs (Schlag et al., 2021).
+Several methods reduce the quadratic complexity of Transformers by making attention linear with respect to context length.
+Linformer projects keys and values into a constant-size matrix, bypassing the scaling with sequence length (Wang et al., 2020).
+Linear Transformer replaces the Softmax function with a decomposable similarity function, achieving linear complexity (Katharopoulos et al., 2020).
+Performer approximates softmax attention using orthogonal random features (Choromanski et al., 2020).
+RetNet adds constant forgetting and rotation (Sun et al., 2023), while Gated Linear Attention introduces learnable forget gates (Yang et al., 2023).
+Linear attention can also be seen as a fast weight network where a slow net adapts a fast network’s parameters online using inputs (Schlag et al., 2021).
 State Space Models
-Instead of trying to linearize transformers, state space models (SSMs) start with parallelizable linear recurrent networks directly.
-Initially, the state transition matrix $A$ is assumed to be constant so that the recurrence can be computed in parallel using a convolution (Li et al., 2022; Gu et al., 2021).
-Subsequent developments include the Diagonal State Space (DSS) model (Gupta et al., 2022), Gated State Space (GSS) models (Mehta et al., 2022), S5 model (Smith et al., 2022), Bidirectional Gated SSM (BiGS)(Wang et al., 2022), H3 model(Fu et al., 2022), and Mamba (Gu & Dao, 2023).
-In addition, there are also works directly trying to make recurrent networks efficient, which often results in a particular form of SSM as well.
-This includes Deep Linear Recurrent Units (LRUs)(Orvieto et al., 2023; De et al., 2024), Hierarchically Gated Linear RNN (HGRN)(Qin et al., 2024b; a), and RWKV (Peng et al., 2023; 2024).
+State space models (SSMs) focus on parallelizable linear recurrent networks.
+Initially, a constant state transition matrix $A$ allows recurrence to be computed via convolution (Li et al., 2022; Gu et al., 2021).
+Key models include Diagonal State Space (DSS) (Gupta et al., 2022), Gated State Space (GSS) (Mehta et al., 2022), S5 (Smith et al., 2022), Bidirectional Gated SSM (BiGS) (Wang et al., 2022), H3 (Fu et al., 2022), and Mamba (Gu & Dao, 2023).
+Efficient recurrent networks often resemble SSMs, such as Deep Linear Recurrent Units (LRUs) (Orvieto et al., 2023; De et al., 2024), Hierarchically Gated Linear RNNs (HGRN) (Qin et al., 2024b; a), and RWKV (Peng et al., 2023; 2024).
 Fast Weight Programmer
-The concept of a network modifying its own weights in response to inputs is not novel and has historical roots in the Fast-weight Programmer (Schmidhuber, 1992; 1993; Schlag & Schmidhuber, 2017; Schlag et al., 2021).
-Specifically, these models propose updating a weight matrix $W\in\mathbb{R}^{d\times m}$ using the outer product of two vectors, expressed as $\Delta W=v(x_{t})\otimes k(x_{t})$.
-This mechanism closely aligns with the principles underlying the Linear Attention model.
-Our framework builds upon the Fast Weight concept by tailoring the weight update process to address a specific online learning objective, thereby extending its applicability and effectiveness in dynamic learning environments.
+The idea of networks modifying their own weights in response to inputs dates back to the Fast-weight Programmer (Schmidhuber, 1992; 1993; Schlag & Schmidhuber, 2017; Schlag et al., 2021).
+These models update a weight matrix $W\in\mathbb{R}^{d\times m}$ via the outer product of two vectors: $\Delta W=x_{t}\otimes k(x_{t})$, a mechanism similar to Linear Attention.
+Our framework extends this concept by adapting the weight update process to suit specific online learning objectives, enhancing its use in dynamic learning environments.
 Concurrent Work
-There are two concurrent works that share similar ideas to ours.
+Two concurrent works share similar ideas with ours.
 Yang et al.
-(2024) propose a chunk-wise parallel form for scaling the DeltaNet (Schlag et al., 2021) to large-scale language modeling tasks (Yang et al., 2024).
-The update rule in DeltaNet can be viewed as conducting the gradient step instead of solving for the closed form of the online regression objective, hence the state transition matrix is $A(x_{t})=(I-\beta_{t}k_{t}k_{t}^{\top})$, which is not guaranteed to have eigenvalues $\leq 1$ (the forgetting can even be negative).
-Therefore, in practice Yang et al.
-(2024) normalize the key vector $k_{t}$ by its $\ell_{2}$ norm (Schlag et al., 2021), which can be restrictive.
-In contrast, the closed-form update in Longhorn ensures that the state transition matrix is always stable.
-In addition, $A(x_{t})$ is a matrix in DeltaNet while we apply a diagonal approximation ($k_{t}^{\odot 2}$).
-Therefore, Longhorn can leverage either parallel scan (as in Mamba) and chunk-wise parallel training (as in GLA), making it as fast as any existing SSM.
-In practice, we provide a parallel scan CUDA kernel, so Longhorn serves as a drop-in replacement for Mamba.
-On the other hand, Sun et al.
-(2024) propose the Test-Time Training framework, where the update of the state is derived from applying one gradient step of an online regression objective (Sun et al., 2024).
-To make the model parallelizable, the authors assume that each gradient step at $x_{t}$ ($t\geq 0$) is conducted at state $s_{0}$, hence deriving a parallel form that utilizes matrix multiplication.
-In contrast, Longhorn solves for the closed-form solution on every token, which potentially leads to a more flexible model.
+(2024) propose a chunk-wise parallel approach to scale DeltaNet (Schlag et al., 2021) for large-scale language modeling.
+DeltaNet’s update rule, viewed as a gradient step for an online regression objective, results in a state transition matrix $A(x_{t})=(I-\beta_{t}k_{t}k_{t}^{\top})$, which can have eigenvalues $>$1, leading to instability.
+To address this, Yang et al.
+(2024) normalize the key vector $k_{t}$ by its $\ell_{2}$ norm, which can be restrictive.
+In contrast, Longhorn ensures stability with a closed-form update, using a diagonal approximation ($k_{t}^{\odot 2}$), allowing for both parallel scan (as in Mamba) and chunk-wise parallel training (as in GLA), making it as fast as existing SSMs.
+Additionally, we provide a parallel scan CUDA kernel, enabling Longhorn to serve as a drop-in replacement for Mamba.
+Sun et al.
+(2024) introduce the Test-Time Training framework, where state updates are derived from a gradient step on an online regression objective.
+To maintain parallelism, they assume each gradient step at $x_{t}$ uses the initial state $s_{0}$, enabling matrix multiplication.
+In contrast, Longhorn computes the closed-form solution for every token, offering greater flexibility.
 5 Experiments
-We conduct experiments on both synthetic benchmarks and the language modeling task to validate the performance of Longhorn:
-  1. 1.
-    To test Longhorn’s performance on memorizing associations, we compare it against other SSMs on the synthetic multi-query associative recall benchmark (Arora et al., 2023).
-We find that Longhorn is the only model that achieves near perfect recall rate when the sequence length scales up to 512 with a hidden dimension of 64).
-  2. 2.
-    To see how Longhorn performs on language modeling tasks, we first conduct the scaling law experiments using the OpenWebText dataset (Gokaslan & Cohen, 2019), with model sizes of 120M and 350M and context length 1024 or 4096, and then compare the validation perplexity of Longhorn against that of the other SSM models.
-We find that Longhorn consistently outperforms baseline deep SSM models across context lengths and model sizes.
-  3. 3.
-    In the end, to see how Longhorn performs on downstream language evaluation tasks, we train a 1.3B language model on the SlimPajama dataset (Soboleva et al., 2023) with 100B tokens, and then compare its downstream evaluation performance across 5 standard benchmarks.
-We find that Longhorn achieves 1.8x speedup sampling efficiency compared to Mamba, and also outperforms GLA and Mamba on average in downstream evaluation tasks.
+We validate Longhorn’s performance through the following experiments:
+1)  We compare Longhorn against other SSMs on the multi-query associative recall benchmark (Arora et al., 2023) and find that Longhorn is the only model to achieve near-perfect recall at sequence lengths up to 512 with a hidden dimension of 64.
+2)  Using the OpenWebText dataset (Gokaslan & Cohen, 2019), we assess Longhorn’s performance on language modeling with model sizes of 120M and 350M, and context lengths of 1024 or 4096, showing it consistently outperforms other SSMs in validation perplexity.
+3)  We train a 1.3B language model on the SlimPajama dataset (Soboleva et al., 2023) with 100B tokens and compare its performance across 8 benchmarks, where Longhorn achieves better final performance and $>$1.8x better sample efficiency than Mamba and GLA.
+4)  We additional apply Longhorn to vision domain and compare it against the Vision Mamba (ViM) (Zhu et al., 2024) model (Appendix 5.5), where Longhorn achieves performance comparable (slightly superior) to that of the ViM model.
 5.1 Multi-Query Associative Recall
-We first consider the synthetic benchmark Multi-Query Associative Recal (MQAR)l (Arora et al., 2023).
-The task is the agent will see a sequence of tokens $\{k_{1},v_{1},k_{2},v_{2},\dots,k_{T},v_{T}\}$, where each consecutive two-tokens become a key-value pair.
-At test time, the agent is provided with multiple $k\sim\{k_{1},\dots k_{T}\}$, the goal is for it to “retrieve” the corresponding values, hence the name multi-query associative recall.
+We first consider the synthetic benchmark Multi-Query Associative Recall (MQAR) (Arora et al., 2023).
+The agent observes a sequence of tokens $\{k_{1},v_{1},k_{2},v_{2},\dots,k_{T},v_{T}\}$, where each consecutive two-tokens become a key-value pair.
+At test time, the agent is provided with multiple $k\sim\{k_{1},\dots k_{T}\}$, the goal is to “retrieve” the corresponding values.
 Following the original benchmark, we consider the sequence length $T\in\{64,128,256,512\}$ and model dimension (size of the latent embedding of a token) $d\in\{64,128,256,512\}$.
@@
 The state-transition matrix is a scalar, 5) BaseConv (Arora et al., 2023), an SSM that combines linear projection with convolution, and 6) Mamba (Gu & Dao, 2023), the state-of-the-art SSM that has data-dependent $A$ and $B$ (See equation 8).
-Each experiment individually searches for the best learning rate from $\{1e-4,4.6e-4,2.2e-3,1e-02\}$ (evenly separated in the log space).
+Each experiment individually searches for the best learning rate from $\{10^{-4},4.6\times 10^{-4},2.2\times 10^{-3},10^{-2}\}$.
 Results are summarized in Figure 4.
-Figure 4: Comparison of Longhorn against state-of-the-art SSM variants on the MQAR benchmark. The y-axis is the recall rate.
+Figure 4: Comparison of Longhorn to state-of-the-art SSMs on the MQAR benchmark. y-axis is the recall rate.
 Observation: From the figure, we can see that Longhorn, which is designed to perform the associative recall task by solving the online prediction objective, outperforms existing SSM variants even at the sequence length of 512 and a small model dimension of 64.
 5.2 Scaling Law on OpenWebText
-In this section we consider language modeling tasks on models with 120M or 350M parameters with 1024 or 4096 context length.
+In this section, we consider language modeling tasks on models with 120M or 350M parameters with 1024 or 4096 context length.
 We choose the OpenWebText dataset as it is small and serves as an easily accessible benchmark for quick benchmarks.
-[^2] We adapt the code from the nanoGPT repo, a minimal reproduction of the GPT-2 model.
-See https://github.com/karpathy/nanoGPT/tree/master.
-Following the Mamba (Gu & Dao, 2023) experiment setting, the training details are provided in Table 2:
-Params n_layers d_model n_heads / d_head Training steps Learning Rate Batch Size Tokens
-125M 12 768 12 / 64 4800 6e-4 0.5M tokens 2.5B
-350M 24 1024 16 / 64 13500 3e-4 0.5M tokens 7B
-Table 2: Training details on OpenWebText.
+[^1] We adapted code from the nanoGPT repository https://github.com/karpathy/nanoGPT, which is a minimal reproduction of GPT-2 model using PyTorch.
+The details about the architecture is provided in Appendix D.
 We consider the following baseline models: LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), Mamba (Gu & Dao, 2023), RWKV (Peng et al., 2023), and GLA (Yang et al., 2023).
-Then we experiment with 1024 or 4096 context length $T$, which model sizes around 120M or 350M.
-Results are summarized in Figure 5 and Table 3.
+Then we experiment with 1024 or 4096 context length $T$ and model sizes around 120M or 350M.
+Results are summarized in Table 1 and Figure 5.
 Figure 5: Scaling law with 1024 and 4096 context length on OpenWebText with various SSM models and the LLaMA (strong Transformer) baseline.
@@
 All models are trained on the OpenWebText dataset (Gokaslan & Cohen, 2019).
-Models vary from around 120M and 350M parameters, with context length 1024 or 4096.
+Models vary from 120-350M parameters and 1024-4096 context length.
 Observation: From the figure and table, we can see that Longhorn consistently outperforms baseline SSMs up to 350M and 4096 context length.
 5.3 Large-scale Language Modeling
-For the large-scale language modeling task, we followed the experiment setting in GLA (Yang et al., 2023), where a model with around 1.3B parameters is trained on the SlimPajama (Soboleva et al., 2023) dataset over 100B tokens with a batch size of 2M tokens.
-We use the AdamW optimizer (Loshchilov & Hutter, 2017) with weight decay of 0.01, and apply cosine learning rate decay with a peak learning rate of $43e-$ 4 and the final learning rate of $53e-$ 5.
-In addition, we use gradient clipping of 1.0.
-Due to the large size of the experiment, we only compare against LLaMA, Mamba and GLA models, all with a context size of 2048.333Our result is in general better than those reported in the GLA paper, this might due to the fact that we are using a different 100B subset of SlimPajama since the original split of the SlimPajama is not provided.
-We evaluate the trained models on a wide range of downstream tasks covering both common-sense reasoning and question-answering.
-In particular, we consider PIQA (Bisk et al., 2020), HellaSwag (Hella) (Zellers et al., 2019), WinoGrande (Wino) (Sakaguchi et al., 2021), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social Interaction QA (SIQA) (Sap et al., 2019), and Boolean questions (BoolQ) (Clark et al., 2019).
-We report the average perplexity across the above eight datasets throughout training in Figure 1 (left). Then we summarize the downstream evaluation results in Table 4.
+For the large-scale language modeling task, we followed the GLA (Yang et al., 2023) setup, training a 1.3B parameter model on the SlimPajama (Soboleva et al., 2023) dataset with 100B tokens and a batch size of 2M.
+We used the AdamW optimizer (Loshchilov & Hutter, 2017) with a weight decay of 0.01, cosine learning rate decay (peak: $43e-$ 4, final: $53e-$ 5), and gradient clipping of 1.0.
+Comparisons were made against LLaMA, Mamba, and GLA models (context size: 2048).
+We evaluated on eight standard downstream tasks, including PIQA (Bisk et al., 2020), HellaSwag (Hella) (Zellers et al., 2019), WinoGrande (Wino) (Sakaguchi et al., 2021), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social Interaction QA (SIQA) (Sap et al., 2019), and Boolean questions (BoolQ) (Clark et al., 2019).
+We report the average perplexity across the above eight datasets throughout training in Figure 1 (left).
+Then we summarize the downstream evaluation results in Table 2.
 Model State Size PIQA Hella Wino. ARC-e ARC-c OBQA SIQA BoolQ Avg.
 acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$
-LLaMA 1024K 55.08 55.36 71.73 59.26 32.19 43.35 45.16 62.13 53.03
-GLA 1024K 55.55 49.10 71.12 58.86 28.11 41.67 44.91 59.21 51.07
-Mamba 32K 54.21 53.61 71.67 61.05 30.15 43.94 44.18 59.22 52.25
-Longhorn 32K 55.78 52.30 71.00 60.63 29.53 43.55 44.68 61.29 52.35
+LLaMA 8M 55.08 55.36 71.73 59.26 32.19 43.35 45.16 62.13 53.03
+GLA 512K 55.55 49.10 71.12 58.86 28.11 41.67 44.91 59.21 51.07
+Mamba 64K 54.21 53.61 71.67 61.05 30.15 43.94 44.18 59.22 52.25
+Longhorn 64K 55.78 52.30 71.00 60.63 29.53 43.55 44.68 61.29 52.35
 Table 2: Language modeling results against LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2023).
@@
 5.4 Ablation on Length Extrapolation
-Moreover, we evaluate how Longhorn extrapolates to a context length longer than 2048 (training context length) at inference time.
+We evaluate how Longhorn extrapolates to a context length longer than 2048 (training context length) at inference time.
 In particular, we pick a disjoint validation set from SlimPajama dataset, rearrange it into batches of sequences of length $T\in\{2048,4096,8192,16384,32768\}$, and then evaluate the pretrained model’s perplexity on those sequences.
@@
 Observation: From the figure, we observe that Longhorn successfully extrapolates to contexts up to 16x longer than those used during training, this contrasts with DeltaNet (Yang et al., 2024), which highlights a limitation in that the model cannot extrapolate to longer contexts.
-Moreover, it is expected that LLaMA, which is a Transformer-based model, fails to extrapolates beyond its training context length.
+In contrast, LLaMA, as a Transformer-based model, fails to extrapolate beyond its training context length.
+5.5 Vision State Space Models
+In addition to language tasks, recent works have also applied state space models to the vision domain, leveraging their superior training efficiency.
+In particular, following the Vision Mamba (ViM) (Zhu et al., 2024), we conduct experiments on the ImageNet (Deng et al., 2009) classification task.
+Similar to ViM, We apply a bi-directional scan with Longhorn SSM (ViL) and compare the results with ViM on both the tiny and small configurations described in the ViM paper.
+Model # Param Top-1 Accuracy
+ViM-Tiny 7M 76.1
+ViL-Tiny (ours) 7M 76.4
+ViM-Small 26M 80.5
+ViL-Small (ours) 26M 80.7
+Table 3: Top-1 Accuracy on ImageNet classification for Vision Mamba (ViM) and Vision Longhorn (ViL).
+Observation: The results from Table 3 demonstrate that the Vision Longhorn model (ViL) achieves comparable (slightly better) performance to the original ViM.
+Note that we use the best hyperparameters for ViM without additional tuning, and ViL does not require two additional parameters for the forward and backward $A$ matrices, as they are computed directly based on the key $k$ vector.
 6 Conclusion and Future Work
-In this work, we introduce a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective.
-We propose a straightforward online regression objective, adopting its implicit closed-form update to define our model, which we refer to as Longhorn.
-Notably, Longhorn is highly parallelizable and achieves state-of-the-art performance in both synthetic sequence modeling and language modeling tasks.
-For future research, an intriguing avenue would be exploring other online learning objectives that can be efficiently implemented on modern hardware.
-Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improves performance.
+This work introduces a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective.
+Based on this, we propose Longhorn, a novel SSM model that explicitly solves online associative recall in closed form.
+Longhorn is parallelizable and achieves state-of-the-art performance among SSMs on MQAR, language modeling, and image classification tasks.
+One future direction is to explore other online learning objectives.
+Additionally, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improves performance.
 We anticipate similar benefits for Longhorn.
+Appendix A Prior Deep State Space Models
+Example A.1 (Linear Attention Variants).
+Linear Attention (LA) (Katharopoulos et al., 2020), Retention Network (RetNet) (Sun et al., 2023), and Gated Linear Attention (GLA) (Yang et al., 2023) all assume $A_{t},B_{t}$ yield rank-1 (or even constant) outputs:
+$S_{t}=A_{t}\odot S_{t-1}+v(x_{t})\otimes k(x_{t}),~{}~{}~{}~{}\text{with}~{}~{}\begin{cases}A_{t}=1&\quad\text{(LA)}\\ A_{t}=c\in[0,1]&\quad\text{(RetNet)}\\ A_{t}=1\otimes\alpha(x_{t})&\quad\text{(GLA)}\end{cases},$
+where $S_{t}\in\mathbb{R}^{d\times m}$, $v(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ are linear mappings of $x_{t}$, and $\otimes$ denote the outer product.
+In practice, one can use $h$ heads as in the multi-head attention to save some computation, where the $m$ and $d$ dimensions are divided into $h$ groups and each group performs its own LA variant.
+The outer product complexity reduces to $\mathcal{O}(h*m/h*d/h=md/h)$.
+But then the effective size of $S_{t}$ also shrinks to $md/h$.
+Example A.2 (Mamba (Gu & Dao, 2023)).
+The Mamba architecture is derived by discretizing a continuous linear dynamics. Its discretized update is:
+$\begin{split}S_{t}&=A_{t}\odot S_{t-1}+B_{t},~{}~{}~{}~{}\text{where}\\ A_{t}&=\exp(A\odot(\varepsilon(x_{t})\otimes 1)),\qquad B_{t}=(\varepsilon(x_{t})\odot x_{t})\otimes k(x_{t}).\end{split}$ (8)
+where $S_{t}\in\mathbb{R}^{d\times m}$ with $m=16$ by default, $\varepsilon(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ linear mappings of $x_{t}$, and $A\in\mathbb{R}^{d\times m}$ is a data independent (not depending on $x_{t}$ trainable weight matrix.
+In Mamba, both $A_{t}$ and $B_{t}$ depend on $\varepsilon(x_{t})$, which represents the step size for the SSM update.
+In practice, Mamba does not use multiple heads as in linear attention variants.
+Perhaps the main reason is that given a fixed $m$ and $d$, the largest memory state will be with $h=1$ (as the effective size of $S_{t}$ is $md/h$).
+In addition, Mamba’s output is $o_{t}=C(x_{t})S_{t}+D_{t}\odot x_{t}$, which has an additional residual part $D_{t}\odot x_{t}$.
+Example A.3 (Griffin (De et al., 2024)).
+In Mamba and the linear attention variants, the outer product serves as a critical role in lifting vectors to matrices.
+The recent Griffin architecture abandons the outer product and performs pure elementwise product:
+$s_{t}=a(x_{t})\odot s_{t-1}+\sqrt{1-a(x_{t})}\odot i(x_{t})\odot x_{t},$
+where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$.
+This yields smaller memory states, but in practice, Griffin is combined with local attention (i.e., the sliding-window self-attention) to strengthen its capability.
+Example A.4 (RWKV (Peng et al., 2023)).
+The original RWKV also performs elementwise recurrence. It maintains a state of ratio form $s_{t}=u_{t}/z_{t}$, where $u_{t},z_{t}$ are updated separately by two SSMs:
+$s_{t}$ $=u_{t}/z_{t}$
+$u_{t}$ $=\exp(-w)\cdot u_{t-1}+\exp(k(x_{t}))\odot v(x_{t}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}z_{t}=\exp(-w)\cdot z_{t-1}+\exp(k(x_{t})),$
+where all the vectors are of size $\mathbb{R}^{d}$, and $w>0$ is a trainable weight for controlling the forgetting.
+In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with the outer product, which makes it more similar to an LA variant.
+Example A.5 (HGRN2 (Qin et al., 2024a)).
+The Gated Linear RNNs with State Expansion (HGRN2) model is represented with the following recurrence:
+$S_{t}=(1\otimes f(x_{t}))\odot S_{t-1}+i(x_{t})\otimes(1-f(x_{t})).$
+Here, $f(x_{t})\in[0,1]$ is the forget gate, $(1-f(x_{t}))$ is the input gate, and $i(x_{t})$ is the input vector. HGRN2 thus resembles an RNN.
+Appendix B Existing State Space Models’ Online Objectives
+We reverse-engineer some existing deep SSMs’ online learning objectives in Table 4.
+Method Online Learning Objective $L_{t}(s)$ (assume $x_{t}\in\mathbb{R}$) Online Update
+LA $\left\lVert S-S_{t-1}\right\rVert^{2}_{\text{F}}-2\langle Sk_{t},x_{t}\rangle$ $S_{t}=S_{t-1}+x_{t}\otimes k_{t}$
+RetNet $\gamma\left\lVert S-S_{t-1}\right\rVert^{2}+(1-\gamma)\left\lVert S\right\rVert^{2}_{\text{F}}-2\langle Sk_{t},x_{t}\rangle$ $S_{t}=\gamma S_{t-1}+x_{t}\otimes k_{t}$
+GLA $\left\lVert S-S_{t-1}\mathrm{diag}(\alpha_{t})\right\rVert^{2}_{\text{F}}+2\langle Sk_{t},x_{t}\rangle$ $S_{t}=S_{t-1}\mathrm{diag}(\alpha_{t})+x_{t}\otimes k_{t}$
+Griffin $\left\lVert\sqrt{\alpha_{t}}\odot(s-s_{t-1})\right\rVert^{2}+\left\lVert\sqrt{1-\alpha_{t}}\odot s\right\rVert^{2}-2\sqrt{1-\alpha_{t}}\odot s\odot i_{t}\odot x_{t}$ $s_{t}=\alpha_{t}\odot s_{t-1}+\sqrt{(1-\alpha_{t})}\odot i_{t}\odot x_{t}$
+Longhorn $\left\lVert S-S_{t-1}\right\rVert^{2}_{\text{F}}+\left\lVert Sk_{t}-x_{t}\right\rVert^{2}_{\mathrm{diag}(\beta_{t})}$ $S_{t}=(1_{m\times n}-\varepsilon_{t}\otimes k_{t}^{\odot 2})\odot S_{t-1}+$
+$(\varepsilon_{t}\odot x_{t})\otimes k_{t},~{}~{}\varepsilon_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$
+Table 4: Some of the existing SSMs and their corresponding online learning objectives/updates.
+Appendix C Proof
+This section provides the proof for Theorem 6.
+Given the Longhorn’s objective $S_{t}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d\times m}}\left\{||S-S_{t-1}||^{2}_{\text{F}}+||Sk_{t}-x_{t}||^{2}_{\mathrm{diag}(\beta_{t})}\right\}$, we have the following theorem:
+Theorem C.1.
+The closed form solution for $S_{t}$ for objective in Equation 5 is
+$S_{t,i}=(I-\varepsilon_{t,i}k_{t}k_{t}^{\top})S_{t-1,i}+\varepsilon_{t,i}k_{t}x_{t,i},~{}~{}~{}\text{where}~{}\varepsilon_{t,i}=\frac{\beta_{t,i}}{1+\beta_{t,i}k_{t}^{\top}k_{t}}\in[0,\infty).$ (9)
+Proof.
+As the objective in equation 5 is in a quadratic form with respect to $s$, there is a unique minimum.
+Observe that each row of $S$ (e.g., $S_{i}$) optimizes the objective independently, therefore we can solve the solution row-wise.
+By setting the derivative of $\nabla_{S_{i}}L_{t}=0$, we have:
+$\nabla_{S_{i}}L_{t}=0$ $\Longleftrightarrow(S_{i}-S_{t-1,i})+\beta_{t,i}(S_{i}^{\top}k_{t}-x_{t,i})k_{t}=0$
+$\Longleftrightarrow(I+\beta_{t,i}k_{t}k_{t}^{\top})S_{i}=S_{t-1,i}+\beta_{t,i}k_{t}x_{t,i}$
+$\underbrace{\Longleftrightarrow}_{(3)}S_{i}=\bigg{(}I-\frac{\beta_{t,i}}{1+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}S_{t-1,i}+\bigg{(}I-\frac{\beta_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}\beta_{t,i}k_{t}x_{t,i}$
+$\Longleftrightarrow\bigg{(}I-\frac{\beta_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}S_{t-1,i}+\frac{(I+\beta_{t,i}k_{t}^{\top}k_{t}-\beta_{t,i}k_{t}k_{t}^{\top})\beta_{t,i}k_{t}x_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}$
+$\underbrace{\Longleftrightarrow}_{(5)}\bigg{(}I-\frac{\beta_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}S_{t-1,i}+\frac{\beta_{t,i}k_{t}x_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}$
+(3) is derived from the fact that $(I+\beta_{t,i}k_{t}k_{t}^{\top})^{-1}=(I-\frac{\beta_{t,i}k_{t}k_{t}^{\top}}{1+\beta_{t,i}k_{t}^{\top}k_{t}})$ by the Sherman–Morrison formula.
+(5) is derived by noticing that $k_{t}^{\top}k_{t}k_{t}x_{t,i}-k_{t}k_{t}^{\top}k_{t}x_{t,i}=0$.
+∎
+Appendix D Additional Experiment Details
+We provide the architecture detail for conducting the scaling law experiments on OpenWebText in Table 5. The architecture configs follow exactly from the Mamba paper (Gu & Dao, 2023).
+Params n_layers d_model n_heads / d_head Training steps Learning Rate Batch Size Tokens
+125M 12 768 12 / 64 4800 6e-4 0.5M tokens 2.5B
+350M 24 1024 16 / 64 13500 3e-4 0.5M tokens 7B
+Table 5: Training details on OpenWebText.